### v2

# Entropy formula for surface diffeomorphisms

Abstract
Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact surface $M$ with $h_{\rm top}(f)\geq\frac{\lambda^{+}(f)}{r}$ where $\lambda^{+}(f):=\lim_{n\to+\infty}\frac{1}{n}\max_{x\in M}\log\left\|Df^{n}_{x}\right\|$. We establish an equivalent formula for the topological entropy:
$h_{\rm top}(f)=\lim_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx.$
We also characterize the topological entropy via the volume growth of curves and several applications are presented. Our approach builds on the key ideas developed in the works of Buzzi-Crovisier-Sarig (Invent. Math., 2022) and Burguet (Ann. Henri Poincaré, 2024) concerning the continuity of the Lyapunov exponents.
MSC Class : 37A05; 37B40; 37C40; 37D25; 37E30
1 Introduction
1.1 Statement of the main results
Let $f$ be a diffeomorphism on a compact surface $M$. Write
$\lambda^{+}(f):=\lim_{n\to+\infty}\frac{1}{n}\max_{x\in M}\log\left\|Df^{n}_{x}\right\|.$
The sequence $\left\{\max_{x\in M}\log\left\|Df^{n}_{x}\right\|\right\}_{n\geq 1}$ is sub-additive, and hence the above limit exists.
For surface diffeomorphisms with large entropy, we establish an equivalent formula for the topological entropy in terms of the volume growth of the tangent cocycle :
Theorem A.
For any $C^{r}$ ($r>1$) diffeomorphism $f$ on a compact surface $M$ with $h_{\rm top}(f)\geq\frac{\lambda^{+}(f)}{r}$, we have
$h_{\rm top}(f)=\lim_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx.$
Let us first make some remarks about Theorem A.
  • Note that the integral is taken with respect to the Lebesgue measure. The Lyapunov regular set (that is, the set of points with well-defined Lyapunov exponents) usually has zero Lebesgue measure and therefore does not directly contribute to the integral above. Even for ergodic volume-preserving Anosov systems-where Lebesgue almost every point has a positive Lyapunov exponent equal to the entropy of the volume measure (Lebesgue measure)-the integral appearing in Theorem A does not converge to the entropy of the volume measure. Instead, it converges to the topological entropy. In this sense, for sufficiently large time $n$, it is the exceptional points (those for which the derivative has not yet grown at the rate of the positive Lyapunov exponent up to time $n$) that dominate the integral. The order of the operations $\log$ and $\int$ is crucial (see Corollary C).
  • In the special case $r=\infty$, no additional term is added to the topological entropy, and our result thus implies the main results of Kozlovski [23] for surface diffeomorphisms. Also note that Theorem A is formulated directly in terms of the derivative $Df^{n}_{x}$, rather than the induced map $(Df^{n}_{x})^{\wedge}$ between exterior algebras of the tangent spaces which is considered by Kozlovski [23]. Although the equivalence of these two quantities for evaluating the topological entropy is not apparent from their respective forms, it is nonetheless natural to work directly with $Df^{n}_{x}$ in the two-dimensional setting, after a more careful analysis. The use of the derivative $Df^{n}_{x}$ in Theorem A might provide a more effective approach for numerically estimating the topological entropy.
    The extension of the formula in Theorem A from the $C^{\infty}$ to the $C^{r}$ setting is far from routine and requires significant improvements in Yomdin theory. The key obstruction lies in the presence of local complexity (indicated by the Yomdin term $\frac{\lambda^{+}(f)}{r}$) . In the $C^{\infty}$ case, this obstruction disappears which drastically simplifies the problem  [^1] This is also the case for Anosov systems, where no local complexity is present, and the formula in Theorem A can be proved in a relatively straightforward way, without any assumption on the topological entropy or the regularity (see, e.g., Kifer-Newhouse [21], Yang-Zang [34]).. In contrast, in the $C^{r}$ setting, the persistence of a nontrivial local complexity term makes the analysis substantially more delicate. We refer to the background in Section 1.2 for a detailed comparison between the two settings.
  • The threshold $\frac{\lambda^{+}(f)}{r}$ for the topological entropy is typically not expected to be attained in Yomdin-type results. For instance:
    • For $C^{r}$ ($r>1$) surface diffeomorphisms with $h_{\rm top}(f)>\frac{\lambda^{+}(f)}{r}$, Buzzi-Crovisier-Sarig [8] (see also subsequent extensions by Buzzi-Luo-Yang [11], Burguet-Luo-Yang [6] and Zang [37]) showed that there are at most finitely many ergodic measures of maximal entropy.
    • For $C^{r}$ ($r>1$) surface diffeomorphisms with $h_{\rm top}(f)>\frac{\lambda^{+}(f)}{r}$, Burguet [4] showed if $\mu_{n}$ is a sequence of ergodic measures with $h(f,\mu_{n})\to h_{\rm top}(f)$ and $\mu_{n}\to\mu$ for some invariant measure $\mu$, then $\mu$ is a measure of maximal entropy and the positive Lyapunov exponent of $\mu_{n}$ converges to that of $\mu$.
    • For $C^{r}$ ($r>1$) surface diffeomorphisms, Burguet [5] also showed that if Lebesgue almost every point has maximal Lyapunov exponent strictly larger than $\frac{\lambda^{+}(f)}{r}$, then the system admits an SRB measure.
    All these results rely on the assumption that certain invariants-such as the topological entropy or Lyapunov exponents-are strictly larger than the threshold $\frac{\lambda^{+}(f)}{r}$. The critical value $\frac{\lambda^{+}(f)}{r}$ itself is delicate and remains poorly understood. In contrast, Theorem A does not suffer from this limitation, as it concerns a formula for the topological entropy itself.
    • The case of systems with low regularity ($r=1$) is not understood by us.
    • The condition $h_{\rm top}(f)\geq\frac{\lambda^{+}(f)}{r}$ appears to be critical. There are examples for interval maps (see Kozlovski [23]) with $h_{\rm top}(f)<\frac{\lambda^{+}(f)}{r}$, in which the topological entropy vanishes while the volume growth of the tangent cocycle appearing in Theorem A is positive. Although this does not immediately yield counterexamples in our setting of surface diffeomorphisms, we do not expect a fundamental gap between the situations of interval maps and surface diffeomorphisms  [^2] We also note that Kozlovski [23, Page 2] mentioned, without providing a precise reference, that Misiurewicz constructed counterexamples for surface diffeomorphisms where the topological entropy is strictly smaller than the volume growth..
    • Recently, S. Ben Ovadia and Burguet [2] studied the Viana conjecture in arbitrary dimensions using a refined high-dimensional version of Yomdin theory. This work may shed light on possible extensions of Theorem A to higher dimensions.
We also provide a formula expressing the topological entropy in terms of the volume growth of the sub-manifolds which extends the classical results in Yomdin theory [35] in $C^{\infty}$ setting.
The volume (i.e., arc length) of a curve $\sigma:[0,1]\to M$ is denoted by ${\rm Vol}\left(\sigma\right)$.
Theorem B.
For any $C^{r}$ ($r>1$) diffeomorphism $f$ on a compact surface $M$ with $h_{\rm top}(f)\geq\frac{\lambda^{+}(f)}{r}$, we have
$h_{\rm top}(f)$ $=\lim_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}\left(f^{n}\left(\sigma\right)\right)$
$=\sup_{\sigma}\limsup_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(\sigma\right)\right)$
$=\sup_{\sigma}\liminf_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(\sigma\right)\right)$
where the supremum is taken over all $C^{r}$ embedded curves $\sigma:[0,1]\to M$ with $\left\|\sigma\right\|_{C^{r}}\leq 1$ and $\left\|d_{t}\sigma\right\|\geq 1/2,\,t\in[0,1]$  [^3] These conditions ensure that the curves are neither degenerate nor highly oscillatory and are imposed only for technical convenience. One may allow more general families..
We make some comments.
  • In general, beyond our $C^{r}$ setting together with the condition $h_{\rm top}(f)\geq\frac{\lambda^{+}(f)}{r}$, it is quite subtle to establish an equivalence between the two kinds of volume growth of the sub-manifolds in Theorem B :
    $\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}\left(f^{n}\left(\sigma\right)\right)\stackrel{{\scriptstyle?}}{{=}}\sup_{\sigma}\limsup_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(\sigma\right)\right).$
    This question was implicitly raised in an unpublished manuscript of Buzzi [7, Remark 1.2].
  • If the topological entropy is strictly larger than $\frac{\lambda^{+}(f)}{r}$, then there exists an invariant measure $\mu$ of maximal entropy ([4, 10]). Moreover, by a result of Cogswell [13] (see also Zang [36]), the supremum in Theorem B is attained by the local unstable manifold $W^{u}_{\rm loc}(x)$ for $\mu$-almost every point $x$. In fact, Corollary E shows that the supremum is attained for any ergodic measure $\mu$ with $h(f,\mu)>\frac{\lambda^{+}(f)}{r}$, and not only for measures of maximal entropy as one might expect.
  • It is worth noting that Theorem B actually implies Theorem A. Indeed, by a Fubini-type argument (see the proof of Theorem 2.2 in Section 5.1), there exists a constant $K>0$ such that for any $n\geq 1$,
    $\int_{M}\left\|Df^{n}_{x}\right\|\,dx\leq K\cdot\sup_{\sigma}{\rm Vol}(f^{n}\sigma).$
    As a consequence, by Theorem B, we have
    $\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\leq\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}(f^{n}\sigma)=h_{\rm top}(f).$
    The reverse inequality ($\geq$) follows from pure Pesin theory (see Theorem 5.2).
    However, we will directly prove Theorem A using a more fundamental maximum bound, namely Theorem 2.2. This approach offers a deeper insight into the mechanism governing the volume growth of tangent cocycle. But there is no essential difference between these two approaches, since both ultimately follow from Theorem 2.1.
1.2 Background and applications
The study of the volume growth of the tangent cocycle:
$\frac{1}{n}\log\int\left\|(Df^{n}_{x})^{\wedge}\right\|\,dx$
appears to go back to Sacksteder-Shub[31]. Here $(Df^{n}_{x})^{\wedge}$ denotes the linear map induced by $Df^{n}_{x}$ on the exterior algebras of the tangent spaces $T_{x}M$ and $T_{f^{n}x}M$, and $\left\|\cdot\right\|$ is the operator norm induced by the Riemannian metric.
Based on Pesin theory, an inequality was obtained by Przytycki [30] for $C^{r}$ ($r>1$) diffeomorphisms and by Newhouse [27] for $C^{r}$ ($r>1$) maps on a compact manifold $M$ of arbitrary dimension:
$\liminf_{n\to+\infty}\frac{1}{n}\log\int\left\|(Df^{n}_{x})^{\wedge}\right\|\,dx\geq h_{\rm top}(f).$ (1)
Subsequently, Kozlovski [23](see also Klapper-L. S. Young [22]) showed that, for $C^{\infty}$ diffeomorphisms, by Yomdin theory [35], a lower bound for the topological entropy is provided, i.e.,
$\limsup_{n\to+\infty}\frac{1}{n}\log\int\left\|(Df^{n}_{x})^{\wedge}\right\|\,dx\leq h_{\rm top}(f).$ (2)
Combining (1) and (2), we obtain an equality for the topological entropy. One might wonder what the $C^{r}$ version of the above formula is for finite $r$. Indeed, in the $C^{r}$ setting, one may expect an inequality that combines the topological entropy with the Yomdin term:
$\limsup_{n\to+\infty}\frac{1}{n}\log\int\left\|(Df^{n}_{x})^{\wedge}\right\|\,dx\leq h_{\rm top}(f)+\frac{\lambda^{+}(f)}{r}.$ (3)
In the $C^{\infty}$ case, the Yomdin term $\frac{\lambda^{+}(f)}{r}$ vanishes, which yields the inequality (2). In contrast, for finite $r$, the inequality (3) cannot be reduced to inequality (2), as the Yomdin term does not vanish in general. Rather than the additive bound of the topological entropy and the Yomdin term appearing in (3), we obtain a maximum bound that yields a substantial improvement (see Theorem 2.2):
$\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\leq\max\left\{h_{\rm top}(f),\,\,\frac{\lambda^{+}(f)}{r}\right\}.$ (4)
If $h_{\rm top}(f)\geq\frac{\lambda^{+}(f)}{r}$, inequality (4) immediately implies inequality (2). Together with inequality (1) which holds for any $C^{r}$ ($r>1$) diffeomorphism, this yields Theorem A.
There are also related results in the partially hyperbolic setting (see the works of Saghin [32], Yang-Zang [34] and Guo-Liao-Sun-Yang [16]), as well as in the setting of random dynamical systems (see Ma [26]).
Next, we present two applications (Corollary C, Corollary D) of Theorem A, as well as two applications (Corollary E, Corollary F) of Theorem B.
The order of the operations $\log$ and $\int$ in Theorem A is crucial in the following sense.
Corollary C.
For any $C^{r}$ ($r>1$) diffeomorphism $f$ on a compact surface $M$ with $h_{\rm top}(f)>\frac{\lambda^{+}(f)}{r}$  [^6] Note that here we have to assume a strict inequality. Otherwise, Burguet’s result [5] does not guarantee the existence of an SRB measure., if
$\limsup_{n\to+\infty}\frac{1}{n}\int_{M}\log\left\|Df^{n}_{x}\right\|\,dx=\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx,$
then there exists finitely many ergodic SRB measures. Moreover, each of them is a measure of maximal entropy  [^7] A measure of maximal entropy is an $f$-invariant probability measure whose metric entropy is equal to the topological entropy $h_{\rm top}(f)$..
Proof.
By Reverse Fatou Lemma (see Lemma 6.4) and Theorem A,
$\int_{M}\limsup_{n\to+\infty}\frac{1}{n}\log\left\|Df^{n}_{x}\right\|\,dx$ $\geq\limsup_{n\to+\infty}\frac{1}{n}\int_{M}\log\left\|Df^{n}_{x}\right\|\,dx$
$=\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx$
$=h_{\rm top}(f)$
$>\frac{\lambda^{+}(f)}{r}.$
We claim that for Lebesgue almost every $x$,
$\limsup_{n\to+\infty}\frac{1}{n}\log\left\|Df^{n}_{x}\right\|=h_{\rm top}(f).$ (5)
Because otherwise, since the integral of the left side is no less than $h_{\rm top}(f)$, there must exist some set $A$ with positive Lebesgue measure such that
$\limsup_{n\to+\infty}\frac{1}{n}\log\left\|Df^{n}_{x}\right\|>h_{\rm top}(f)>\frac{\lambda^{+}(f)}{r},\quad\forall x\in A.$
By Burguet’s result [5], there is an ergodic SRB measure $\mu$ whose positive Lyapunov exponent is strictly larger than $h_{\rm top}(f)$. This is a contradiction since the positive exponent of an ergodic SRB measure coincides with its entropy.
Again, by Burguet’s result [5], under condition (5), there are at most countably many ergodic SRB measures $\{\mu_{i}\}_{i\in\mathbb{N}}$ such that Lebesgue almost every $x$ lies in the basin of some $\mu_{i}$ and each $\mu_{i}$ carries maximal entropy. Indeed, in our setting there are only finitely many such ergodic SRB measures, as a direct consequence of the main results of Buzzi-Crovisier-Sarig [8], which assert that there are at most finitely many ergodic measures of maximal entropy.
The following result, while not new  [^9] Although this result does not appear in this exact form, Corollary D can also be deduced from Burguet’s result [5] by considering SRB measures, in a way similar to the proof of Corollary C. Here we give a technically lighter proof using Theorem A., is another interesting consequence of Theorem A. Denote the usual maximal Lyapunov exponent of $x$ by
$\lambda^{+}(f,x):=\limsup_{n\to+\infty}\frac{1}{n}\log\left\|Df^{n}_{x}\right\|.$
Corollary D.
For any $C^{r}$ ($r>1$) diffeomorphism $f$ on a compact surface $M$ with $h_{\rm top}(f)\geq\frac{\lambda^{+}(f)}{r}$,
$\lambda^{+}(f,x)\leq h_{\rm top}(f)$
for Lebesgue almost every point $x$.
Proof.
By a general result in measure theory (see Lemma 6.5) and Theorem A,
$\lambda^{+}(f,x)\leq\lim_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx=h_{\rm top}(f)$
for Lebesgue almost every point $x$. ∎
The following Corollay E is an application of Theorem B, showing that the maximal volume growth is attained w.r.t. any ergodic measure $\mu$ satisfying $h(f,\mu)>\frac{\lambda^{+}(f)}{r}$, rather than being restricted to measures of maximal entropy. This conclusion is somewhat unexpected. For instance, consider an SRB measure $\mu$: although its unstable manifolds carry $\mu$-typical points with full induced Lebesgue measure, their volume growth does not converge to the positive Lyapunov exponent (which coincides with the entropy of $\mu$), but instead to the topological entropy.
It is instructive to compare this phenomenon with the so-called Almost Anosov systems constructed by Hu-Young [17], where Lebesgue almost every point has zero Lyapunov exponents, yet a measure of maximal entropy still exists. In particular, a curve along which Lebesgue almost every point has zero Lyapunov exponents may nevertheless exhibit maximal volume growth, equal to the topological entropy. On the other hand, it is not hard to see: a curve along which Lebesgue almost every point has positive Lyapunov exponents cannot have zero volume growth.
The proof of Corollary E relies on background from homoclinic relations (see [8]) and Pesin theory (see [1]). Here we only provide an outline of the argument, omitting the technical details.
Corollary E.
For any $C^{r}$ ($r>1$) transitive diffeomorphism $f$ on a compact surface $M$ and any ergodic measure $\mu$ with $h(f,\mu)>\frac{\lambda^{+}(f)}{r}$, we have for $\mu$-almost every $x$,
$h_{\rm top}(f)=\lim_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(W^{u}_{\rm loc}(x)\right)\right)$
where $W^{u}_{\rm loc}(x)$ denotes the local unstable manifold at $x$.
Outline of the proof.
Since $f$ is transitive, by [8, Theorem 6.12], there is a unique homoclinic class supports all ergodic measures $\mu$ with $h(f,\mu)>\frac{\lambda^{+}(f)}{r}$. Moreover, this homoclinic class supports a unique ergodic measure $\nu$ of maximal entropy  [^10] The existence of measures of maximal entropy for the $C^{r}$ case is established by Burguet [4, Corollary 1] and the uniqueness is established by Buzzi-Crovisier-Sarig [8, Main Theorem Revisited].. Hence $\mu$ must be homoclinically related to $\nu$, i.e., there are two subsets $R_{\mu},R_{\nu}$ with $\mu(R_{\mu}),\nu(R_{\nu})>0$ such that for any $x\in R_{\mu},y\in R_{\nu}$, the unstable manifold $W^{u}(x)$ intersects the stable manifold $W^{s}(y)$ transversely at some points, and likewise $W^{s}(x)$ intersects $W^{u}(y)$ transversely at some points. Note that for $\nu$-almost every $y\in R_{\nu}$,
$\nu_{y}\left(R_{\nu}\right)>0$
where $\nu_{y}$ denotes the conditional measure  [^11] In fact, conditional measures are defined with respect to subordinate partitions and we disregard this point here. See Ledrappier-Young [18] for further details. of $\nu$ on $W^{u}_{\rm loc}(y)$. We fix any such $y\in R_{\nu}$. Let $z$ be a transverse point of $W^{u}(x)$ and $W^{s}(y)$. Let $m_{1}\in\mathbb{N}$ be large enough such that
$z\in f^{m_{1}}\left(W^{u}_{\rm loc}\left(f^{-m_{1}}\left(x\right)\right)\right).$
By Inclination Lemma  [^12] The Inclination Lemma [8, Lemma 2.7] requires $y$ to be recurrent. As $\nu$-almost every point in $R_{\nu}$ is recurrent (see [8, Section 2.3(e),(f)]), we may assume without loss of generality that $y$ is recurrent., there are $m_{2}\in\mathbb{N}$ and some curve $D$ with
$z\in f^{-m_{2}}\left(D\right)\subset f^{m_{1}}\left(W^{u}_{\rm loc}\left(f^{-m_{1}}\left(x\right)\right)\right)$
such that $D$ is sufficiently $C^{1}$-close to $W^{u}_{\rm loc}(y)$. Consequently, the stable holonomy map $\pi_{s}:\Lambda\to D$ is well defined on some Pesin set  [^13] Here $\Lambda$ refers to the standard Pesin set, where points admit stable and unstable manifolds with uniform size, and the angle between the stable and unstable directions is uniformly bounded away from zero. See [1] for details on Pesin theory. $\Lambda\subset W^{u}_{\rm loc}(y)$ with $\nu_{y}(\Lambda)>0$. See Figure 1.
Figure 1: The stable holonomy map $\pi_{s}$
Let us first briefly recall the argument in [36] (see also [13], [27] for similar ideas) establishing
$h(f,\nu)\leq\liminf_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(W^{u}_{\rm loc}(y)\right)\right).$ (6)
The fact that $\nu_{y}(\Lambda)>0$ is what ensures sufficient volume growth. Given any small $\varepsilon>0$ and any large $n\in\mathbb{N}$, since $\nu_{y}\left(\Lambda\right)>0$, by a fiber-version of Katok’s entropy formula (see [36, Theorem A (1)]), there is a maximal $(n,\varepsilon)$-separated subset $\Omega_{n,\varepsilon}$ of $\Lambda$ with
$\Omega_{n,\varepsilon}\gtrsim e^{n\cdot\left(h(f,\nu)-\varepsilon\right)}$
such that for any $w\in\Omega_{n,\varepsilon}$,
${\rm Vol}\left(f^{n}\left(B^{u}(w,n,\varepsilon)\right)\right)\gtrsim e^{-n\varepsilon}$
where $B^{u}(w,n,\varepsilon)$ denotes the unstable dynamical ball,
$B^{u}(w,n,\varepsilon):=\left\{w^{\prime}\in W^{u}_{\rm loc}(w):\ d\left(f^{k}(w^{\prime}),f^{k}(w)\right)<\varepsilon\ \text{ for all }0\leq k<n\right\}.$
As a consequence,
${\rm Vol}\left(f^{n}\left(W^{u}_{\rm loc}(y)\right)\right)\gtrsim\#\Omega_{n,\varepsilon}\cdot\min_{w\in\Omega_{n,\varepsilon}}{\rm Vol}\left(f^{n}\left(B^{u}(w,n,\varepsilon)\right)\right)\gtrsim e^{n\cdot\left(h(f,\nu)-2\varepsilon\right)}$
which proves (6) by the arbitrariness of $\varepsilon$. The key observation is that the same argument applies with $D$ in place of $W^{u}_{\rm loc}(y)$. Indeed, since $D$ is connected to $W^{u}_{\rm loc}(y)$ through the stable holonomy on $\Lambda$, it follows that, for any large $n$ and any $w\in\Lambda$, the point $f^{n}\left(\pi_{s}(w)\right)$ remains inside the Pesin chart centered at $f^{n}(w)$. As a consequence,
${\rm Vol}\left(f^{n}\left(B^{u}\left(\pi_{s}(w),n,\varepsilon\right)\right)\right)\approx{\rm Vol}\left(f^{n}\left(B^{u}(w,n,\varepsilon)\right)\right)\gtrsim e^{-n\varepsilon}.$
Hence
${\rm Vol}\left(f^{n}\left(D\right)\right)\gtrsim\#\Omega_{n,\varepsilon}\cdot\min_{w\in\Omega_{n,\varepsilon}}{\rm Vol}\left(f^{n}\left(B^{u}\left(\pi_{s}(w),n,\varepsilon\right)\right)\right)\gtrsim e^{n\cdot\left(h(f,\nu)-2\varepsilon\right)}.$
This shows for any $x\in R_{\mu}$,
$h_{\rm top}(f)=h(f,\nu)\leq\liminf_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(D\right)\right)\leq\liminf_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(W^{u}_{\rm loc}\left(f^{-m_{1}}\left(x\right)\right)\right)\right).$
By Theorem B  [^14] Any local unstable manifold is contained in a curve from the family in Theorem B, as local unstable manifolds are nearly straight at small scales.,
$h_{\rm top}(f)=\lim_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(W^{u}_{\rm loc}\left(f^{-m_{1}}\left(x\right)\right)\right)\right).$
By ergodicity and the backward contraction of the local unstable manifolds, this equality above in fact holds for $\mu$-almost every $x$ without passing to backward iterates (see Lemma 6.6):
$h_{\rm top}(f)=\lim_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(W^{u}_{\rm loc}(x)\right)\right).$
The proof of Corollary E in fact yields the following more general statement: if two hyperbolic ergodic measures $\mu$ and $\nu$ are homoclinically related, then for $\mu$-almost every point $x$,
$h(f,\nu)\leq\liminf_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(W^{u}_{\rm loc}\left(x\right)\right)\right).$
As a consequence of variational principle, for any hyperbolic ergodic measure $\mu$ supported on a homoclinic class $H(\mathcal{O})$ and for $\mu$-almost every $x$,
$h_{\rm top}\left(f,H(\mathcal{O})\right)\leq\liminf_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(W^{u}_{\rm loc}(x)\right)\right)$
where $h_{\rm top}\left(f,H(\mathcal{O})\right)$ denotes the topological entropy of the homoclinic class.
In general, the topological entropy of a curve is bounded above by its volume growth (see Lemma 6.2), and the inequality can be strict. Indeed, topological entropy only reflects the number of dynamical balls, capturing the global complexity, whereas volume growth depends not only on this number but also on the geometric distortion within each dynamical ball, such as folding and bending, which represent local complexity. However, the following corollary shows that when a curve carries full topological entropy, its volume growth attains the maximal rate which coincides with the topological entropy. In this case, the contribution of local complexity inside dynamical balls becomes negligible.
Write $\log^{+}x:=\max\{\log x,\,0\}$.
Corollary F.
[…]
$h_{\rm top}(f)=\limsup_{n\to+\infty}\frac{1}{n}\log^{+}{\rm Vol}\left(f^{n}\left(\sigma\right)\right).$
Proof.
By Lemma 6.2, we have
$h_{\rm top}(f)=h_{\rm top}\left(f,\sigma\right)\leq\limsup_{n\to+\infty}\frac{1}{n}\log^{+}{\rm Vol}\left(f^{n}\left(\sigma\right)\right).$
On the other hand, by Theorem B, the right-hand side is bounded above by $h_{\rm top}(f)$. Therefore, equality holds. ∎
1.3 Some further questions
We believe that the ideas of Theorem A and Theorem B might extend to other topological invariants (see Llibre-Saghin [24] for a survey of various invariants measuring dynamical complexity), such as the growth rate of hyperbolic periodic points. This has been established by Burguet [3] in the $C^{\infty}$ setting and it was conjectured by Burguet in the $C^{r}$ case:
Question 1 ([3]).
For any $C^{r}$ ($r>1$) diffeomorphism $f$ on a compact surface $M$ with $h_{\rm top}(f)>\frac{\lambda^{+}(f)}{r}$ and any $0<\delta<h_{\rm top}(f)$,
$h_{\rm top}(f)=\limsup_{n\to+\infty}\frac{1}{n}\log\#\mathcal{P}^{n}_{\delta}$
where $\mathcal{P}^{n}_{\delta}$ is the set of $n$-periodic points with Lyapunov exponents $\delta$-away from zero.
We also propose some questions related to the formula in Theorem A:
$h_{\rm top}(f)=\lim_{n\to+\infty}a_{n}:=\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx.$
Question 2.
In general, does $a_{n}$ always converge ? If so, what is its limit if it is not the topological entropy ?
Remark.
Note that $a_{n}$ might NOT be a sub-additive sequence, even when $f$ preserves the volume (i.e., when the Lebesgue measure is an invariant measure).
We believe that a deeper formula underlies Theorem A.
Question 3.
Is it reasonable to ask about the rate of convergence in Theorem A ?
Remark.
Note that the definition of $a_{n}$ already contains the factor $\frac{1}{n}\log$”. It might therefore be too ambitious to expect a specific rate of convergence (for instance, exponential). However, we do not know this convincingly. We also note that Buzzi-Crovisier-Sarig [10] have recently shown that measures of maximal entropy are exponentially mixing.
Denote by ${\rm Diff}^{r}(M)$ the space of all $C^{r}$ diffeomorphisms on $M$ equipped with the $C^{r}$ topology. Define the functions $\varphi_{n},\varphi:{\rm Diff}^{r}(M)\to\mathbb{R}$:
$\varphi_{n}\left(g\right):=\frac{1}{n}\log\int_{M}\left\|Dg^{n}_{x}\right\|\,dx,\qquad\varphi:=\limsup_{n\to+\infty}\varphi_{n}.$
Question 4.
Is $\varphi$ upper semi-continuous ?
Remark.
[…]
Question 5.
Does an analogue of Theorem A hold for interval maps ?
Remark.
This is known to hold for $C^{\infty}$ interval maps by the work of Kozlovski [23]. It remains unclear whether new obstructions arise when extending it to $C^{r}$ interval maps. In this paper, we only consider diffeomorphisms.
It is also worth noting that there have been several recent works on the continuity of Lyapunov exponents and the existence of SRB measures for interval maps (see, for example, Delplanque [14], Delplanque-Li [15], and Li [19]).
2 Maximum bounds for volume growth
In this section, we present the main results used to prove Theorem A and Theorem B.
2.1 Mechanism behind the maximum bounds for volume growth
Theorem 2.1.
Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact surface $M$. We have
$\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}(f^{n}\sigma)\leq\max\left\{h_{\rm top}(f),\,\,\frac{\lambda^{+}(f)}{r}\right\}$
where the supremum is taken over all $C^{r}$ embedded curves $\sigma:[0,1]\to M$ with $\left\|\sigma\right\|_{C^{r}}\leq 1$ and $\left\|d_{t}\sigma\right\|\geq 1/2,\,t\in[0,1]$.
Remark.
This maximum bound improves the classical estimate in Yomdin theory [35], where only the following sum bound is obtained:
$\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}(f^{n}\sigma)\leq h_{\rm top}(f)+\frac{\lambda^{+}(f)}{r}.$
Theorem 2.2.
Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact surface $M$. We have
$\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\leq\max\left\{h_{\rm top}(f),\,\,\frac{\lambda^{+}(f)}{r}\right\}.$
The inequalities in Theorems 2.1 and 2.2 can be equivalently expressed in a convex combination form: there is some $\alpha\in[0,1]$ such that
$\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}(f^{n}\sigma)\leq\alpha\cdot h_{\rm top}(f)+(1-\alpha)\cdot\frac{\lambda^{+}(f)}{r},$ (7)
$\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\leq\alpha\cdot h_{\rm top}(f)+(1-\alpha)\cdot\frac{\lambda^{+}(f)}{r}.$
The reason for presenting this convex-combination form is that our proof actually yields a more refined upper bound, in which the coefficient $\alpha\in[0,1]$ reflects how the volume growth is distributed between the entropic contribution and the Lyapunov-type contribution (the Yomdin term). In particular, the argument implicitly determines an admissible choice of $\alpha$ and thus contains strictly more information than the crude bound given by the maximum. However, making this choice explicit would require introducing a substantial amount of additional technical machinery. For this reason, we do not pursue an explicit description of $\alpha$ here in the statements. As a consequence, the statements of the theorems may be equivalently weakened to the simpler form involving $\max\{h_{\rm top}(f),\lambda^{+}(f)/r\}$.
These convex-type upper bounds originate from the idea that dynamical invariants such as volume growth measure complexity from two different perspectives: a global one and a local one. To illustrate this idea, let us consider the volume growth of a curve $\sigma$ and write
${\rm Vol}(f,\sigma):=\limsup_{n\to+\infty}\frac{1}{n}\log{\rm Vol}(f^{n}\sigma).$
At time $n$ and at scale $\varepsilon$, the volume of $f^{n}(\sigma)$ can be estimated by decomposing it into two factors:
[…]
The global complexity is controlled by the topological entropy, while the classical Yomdin theory [35] shows that the local complexity grows at most at the rate $\lambda^{+}(f)/r$. Combining these two estimates yields the classical Yomdin bound:
${\rm Vol}(f,\sigma)\leq h_{\rm top}(f)+\frac{\lambda^{+}(f)}{r}.$ (8)
However, this estimate is in general quite loose, since it implicitly assumes that the global and the local complexities are simultaneously present at all times.
The key insight behind the convex bound in (7) is that these two sources of complexity need not act at the same time. Instead, the time interval $[0,n)$ can be decomposed into two major types of periods:
  • Geometric times (with proportion $\alpha$): during which the curve stretches across the manifold in a globally expanding and essentially straight manner;
  • Neutral times (with proportion $1-\alpha$): during which the curve remains confined inside a single dynamical ball and develops complexity only through local folding.
Figure 2: Growth at geometric and neutral times
During geometric times, one only needs to count the number of dynamical balls that the curve visits, which contributes an entropy-type growth, while no significant local complexity is generated inside each dynamical ball. In contrast, during neutral times, the curve stays inside a single dynamical ball, so the global complexity vanishes, and the growth is entirely governed by local effects, captured by the Yomdin term $\lambda^{+}(f)/r$. See Figure 2.
This separation of time scales explains the convex structure of the resulting bound and clarifies the mechanism responsible for the improvement over the classical Yomdin estimate. As a consequence, the classical Yomdin bound (8) can be refined to the following inequality:
${\rm Vol}(f,\sigma)\leq\alpha\cdot h_{\rm top}(f)+(1-\alpha)\cdot\frac{\lambda^{+}(f)}{r}.$
The separation of time into geometric and neutral components was first introduced by Buzzi-Crovisier-Sarig [9] and was further developed and systematized by Burguet [4]. This idea has since been deepened and applied in several related contexts; see, for instance, [6, 11, 25].
We emphasize that all these works are formulated in the presence of a reference invariant measure, and the quantities under consideration are measure-theoretic invariants, such as metric entropy, Lyapunov exponents. In contrast, our approach focuses on purely topological quantities, namely the volume growth of the sub-manifolds and of the tangent cocycle. The absence of a reference measure leads to several substantial technical difficulties. We mention here two of the main ones.
  • The presence of periodic sources  [^17] A periodic point $p$ is called a periodic source if both Lyapunov exponents of $p$ are positive.. When a reference hyperbolic measure $\mu$ is given, the analysis may be restricted to $\mu$-typical points, which are not influenced by periodic sources. In our setting, however, when considering, for example, the volume growth of a curve $\sigma$, a large portion of points in $\sigma$ may spend a long time near periodic sources. During such times, which still correspond to geometric times, the curve needs time to escape from the ”trapping region” of these sources. As a consequence, one must carefully control the amount of topological entropy produced during these geometric times, avoiding an unnecessary decomposition of the curve into exponentially many pieces before it has moved away from periodic sources. To address this issue, we introduce a finer decomposition of time and define the notion of trapping times, which allows us to prevent an excessive creation of entropy (see Proposition 4.8).
  • Lack of uniform control on the local complexity at geometric times Without a reference measure, it becomes more delicate to justify that no local complexity is generated during geometric times in a uniform way. We overcome this difficulty by exploiting several partitions of the curve, as introduced in Proposition 4.10, together with an upper semi-continuity property for sub-additive sequences, which yields the required control.
2.2 Proofs of Theorem A and Theorem B
We now directly apply the two bounds established above in Theorem 2.1 and 2.2 to prove Theorem A and Theorem B, in order to provide the reader with a clear overview of the global framework.
Proof of Theorem A.
The inequality
$h_{\rm top}(f)\leq\liminf_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx$
is, in a certain sense, already well known (see Przytycki [30], Kozlovski [23]). Its proof relies only on the $C^{r}$ ($r>1$) regularity and follows as a consequence of Pesin theory. The only distinction from the classical formulations is that we work directly with the derivative $Df^{n}_{x}$, rather than with the induced map $(Df^{n}_{x})^{\wedge}$ acting on the exterior algebras of the tangent spaces. This difference is natural and expected in our setting, since we restrict ourselves to dimension two. For the reader’s convenience and completeness, we reproduce this part of the argument in Theorem 5.2.
The reverse inequality is provided by Theorem 2.2:
$\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\leq\max\left\{h_{\rm top}(f),\,\,\frac{\lambda^{+}(f)}{r}\right\}=h_{\rm top}(f).$
The proof of Theorem A is complete. ∎
Proof of Theorem B.
In the following arguments, the notation $\sup_{\sigma}$ refers to the supremum taken over all $C^{r}$ embedded curves $\sigma:[0,1]\to M$ with $\left\|\sigma\right\|_{C^{r}}\leq 1$ and $\left\|d_{t}\sigma\right\|\geq 1/2,\,t\in[0,1]$.
By the work of Cogswell [13] (see also Zang [36]), based purely on Pesin theory, for any hyperbolic ergodic measure $\mu$ and for $\mu$-almost every point $x$,
$h(f,\mu)\leq\liminf_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(W^{u}_{\rm loc}(x)\right)\right)\leq\sup_{\sigma}\liminf_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(\sigma\right)\right)$
where $W^{u}_{\rm loc}(x)$ denotes the local unstable manifold at $x$.
Hence, by the variational principle,
$h_{\rm top}(f)\leq\sup_{\sigma}\liminf_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(\sigma\right)\right)\leq\liminf_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}\left(f^{n}(\sigma)\right).$
The reverse inequality is provided by Theorem 2.1,
$\sup_{\sigma}\limsup_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(\sigma\right)\right)$ $\leq\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}\left(f^{n}(\sigma)\right)$
$\leq\max\left\{h_{\rm top}(f),\,\,\frac{\lambda^{+}(f)}{r}\right\}$
$=h_{\rm top}(f).$
The proof of Theorem B is complete.
It remains to prove Theorem 2.1 (Section 4) and Theorem 2.2 (Section 5).
3 Preliminaries
In this section, we collect several notions and basic facts that will be used throughout the paper. We briefly recall Lyapunov exponents and their associated invariant splittings, and then introduce the induced dynamics on the projective tangent bundle, which provides a convenient framework for describing the measure-theoretic invariants.
3.1 Lyapunov exponents
The topological entropy of $f$ is denoted by $h_{\rm top}(f)$ and the metric entropy of an invariant measure $\mu$ is denoted by $h(f,\mu)$. By Oseledets Multiplicative Ergodic Theorem, there is an invariant set $\mathcal{R}$ (called the Lyapunov regular set) with total measure (i.e., $\mu(\mathcal{R})=1$ for any invariant measure $\mu$) such that for any $x\in\mathcal{R}$, there are a splitting (called the Oseledets splitting) $T_{x}M=E^{1}\oplus E^{2}\oplus\cdots\oplus E^{l}$ and finitely many numbers (called the Lyapunov exponents) $\lambda_{1}(x)>\lambda_{2}(x)>\cdots>\lambda_{l}(x)$ such that for any nonzero vector $v\in E^{j}$, we have
$\lim_{n\to\pm\infty}\frac{1}{n}\log\left\|Df^{n}_{x}(v)\right\|=\lambda_{j}.$
In our special setting (surface diffeomorphism), given an invariant measure $\mu$, there are at most two well defined Lyapunov exponents for $\mu$-almost every $x$. The larger Lyapunov exponent $\lambda_{1}(x)$ is denoted by $\lambda^{+}(f,x)$ and it can be also defined as
$\lambda^{+}(f,x)=\lim_{n\to+\infty}\frac{1}{n}\log\left\|Df^{n}_{x}\right\|.$
Similarly, the smaller Lyapunov exponent is denoted by $\lambda^{-}(f,x)$. We rewrite the Oseledets splitting at $x$ as
$T_{x}M=E^{+}_{x}\oplus E_{x}^{-}.$
An invariant measure $\mu$ is called hyperbolic if for $\mu$-a.e. $x$, one Lyapunov exponents of $x$ is positive and the other is negative. For an ergodic measure $\mu$ with $h(f,\mu)>0$, $\mu$ is hyperbolic by Ruelle’s inequality. We define
$\lambda^{+}(f,\mu):=\int\lambda^{+}(f,x)d\,\mu(x),\quad\lambda^{-}(f,\mu):=\int\lambda^{-}(f,x)d\,\mu(x).$
If $\mu$ is ergodic, then $\lambda^{\pm}(f,\mu)=\lambda^{\pm}(f,x)$ for $\mu$-a.e. $x$.
3.2 Tangent dynamics
With the Riemannian structure inherited from the tangent bundle $TM$, we denote the projective tangent bundle of $M$ by
[…]
Let $\pi:\hat{M}\to M$ be the natural projection. Let $\hat{f}:\hat{M}\to\hat{M}$ be the induced map of $f$ (also called the canonical lift) defined by
$\hat{f}(x,[v]):=\left(f(x),[Df_{x}(v)]\right).$
If $f$ is of class $C^{r}$, then $\hat{f}$ is of class $C^{r-1}$.
Let $\rho:\hat{M}\to\mathbb{R}$ be the continuous function defined by
$\rho(x,[v]):=\log\left\|Df_{x}\left(v\right)\right\|.$
Given an $\hat{f}-$invariant measure $\hat{\mu}$, the projection of $\hat{\mu}$ is defined by $\mu:=\hat{\mu}\circ\pi^{-1}$ and $\hat{\mu}$ is called the lift of $\mu$. $\hat{\mu}$ is called an unstable lift of $\mu$ if for $\hat{\mu}-$almost every point $\hat{x}$,
$\lim_{n\to+\infty}\frac{1}{n}\sum_{i=0}^{n-1}\rho\left(\hat{f}^{i}\left(\hat{x}\right)\right)>0.$ (9)
We define
$\lambda(\hat{f},\hat{\mu}):=\int_{\hat{M}}\rho d\,\hat{\mu}$.
We list some basic properties.
  • Given an $\hat{f}-$invariant measure $\hat{\mu}$, if (9) holds for $\hat{\mu}-$almost every point $\hat{x}$, then the Lyapunov exponent $\lambda^{+}(f,x)$ is positive for $\mu$-almost every $x$, where $\mu$ is the projection of $\hat{\mu}$. This follows from the fact that for $\hat{\mu}-$almost every point $\hat{x}=(x,[v])\in\hat{M}$ with $\|v\|=1$,
    $\frac{1}{n}\sum_{i=0}^{n-1}\rho\left(\hat{f}^{i}(\hat{x})\right)=\frac{1}{n}\log\left\|Df_{x}^{n}(v)\right\|.$
[…]
  • Let $\mu$ be an $f$-invariant measure with $\lambda^{+}(f,x)>0\geq\lambda^{-}(f,x)$ for $\mu$-almost every $x$. Then $\mu$ has a unique unstable lift $\hat{\mu}^{+}$ which can be defined by:
    $\hat{\mu}^{+}:=\int_{\hat{M}}\delta_{(x,E^{+}_{x})}d\,\mu(x).$
    Moreover (refer to [10, Proposition 3.8],
    $\lambda^{+}(f,\mu)=\lambda(\hat{f},\hat{\mu})\quad\text{if and only if}\quad\hat{\mu}=\hat{\mu}^{+}.$
    For more detailed discussions on the lifted dynamics, we refer to [9, Section 3.4] and the references therein.
3.3 Notations and conventions
  • Whenever we say that $\sigma:[0,1]\to M$ is a $C^{r}$ curve, we always assume that it is an embedded curve.
  • We identify $\sigma$ with its image whenever there is no ambiguity; in particular, we freely write $x\in\sigma$.
  • Once a curve $\sigma$ is fixed, for any $x\in\sigma$, we denote by $\hat{x}=(x,[v])\in\hat{M}$ where $v$ is the unit tangent vector of $\sigma$ at $x$.
  • Given $\hat{x}:=(x,[v])\in\hat{M}$ where $v$ is a unit vector in $T_{x}M$, we shall slightly abuse notation and write
    $Df(\hat{x}):=Df_{x}(v)\in T_{f(x)}M.$
  • Write
    $\left\|Df\right\|:=\max_{x\in M}\left\|Df_{x}\right\|,\quad\left\|Df^{-1}\right\|:=\max_{x\in M}\left\|Df^{-1}_{x}\right\|.$
    Note that $\left\|Df\right\|,\left\|Df^{-1}\right\|\geq 1$.
  • Throughout the paper, the notation $\sup_{\sigma}$ refers to the supremum taken over all $C^{r}$ embedded curves $\sigma:[0,1]\to M$ with $\left\|\sigma\right\|_{C^{r}}\leq 1$ and $\left\|d_{t}\sigma\right\|\geq 1/2,\,t\in[0,1]$.
4 Volume growth of the sub-manifolds
The purpose of this section is to establish Theorem 2.1 via two key technical results: Proposition 4.8 and Proposition 4.10.
  • Proposition 4.8 separates the volume growth into the entropic contribution and the Lyapunov-type contribution (the Yomdin term) via Burguet’s reparametrizations, a variant of Yomdin’s reparametrizations.
  • Proposition 4.10 refines Proposition 4.8 by controlling the error term arising from entropic contributions, more precisely, the local volume growth at geometric times in Proposition 4.8. It shows that this error term is negligible in the asymptotic regime.
4.1 Reparametrization
Let $\sigma:[0,1]\to M$ be a $C^{r}$ curve. For an integer $s\geq 1$, the quantity $\left\|d^{s}\sigma\right\|$ denotes the supremum norm of the $s$-th derivative of $\sigma$, namely
$\left\|d^{s}\sigma\right\|:=\sup_{t\in[0,1]}\left\|d^{s}_{t}\sigma\right\|,$
where the norm is taken with respect to the Riemannian metric on $M$. We define the $C^{r}$ norm by
$\left\|\sigma\right\|_{C^{r}}:=\max_{1\leq s\leq r}\,\left\|d^{s}\sigma\right\|.$
We say that $\sigma$ is bounded if
$\max_{2\leq s\leq r}\,\left\|d^{s}\sigma\right\|\leq\frac{1}{6}\cdot\left\|d\sigma\right\|.$
Note, consequently, we have
$\left\|d\sigma\right\|\leq 2\cdot\left\|d_{t}\sigma\right\|,\,\forall t\in[0,1].$ (10)
Moreover, the curve $\sigma$ is said to be $\varepsilon$-bounded if
$\left\|d\sigma\right\|\leq\varepsilon.$
Roughly speaking, a bounded curve is one that is nearly straight, whereas an $\varepsilon$-bounded curve is nearly straight and, in addition, has small length.
We often call an affine map from $[0,1]$ to itself a reparametrization.
The following proposition is a reparametrization result in the spirit of Yomdin theory, developed by Burguet [4, Lemma 12]. It shows that an $\varepsilon$-bounded curve can be decomposed, via affine reparametrizations, into finitely many pieces on which the image under $f$ remains bounded. Moreover, the number of such pieces and the sizes of the reparametrizations are quantitatively controlled by the expansion rates along the curve.
The ceiling function $\left\lceil x\right\rceil$ assigns to each real number $x$ the smallest integer $n$ such that $n\geq x$. Recall that, given a curve $\sigma$, $\hat{x}:=(x,[v])\in\hat{M}$ where $v$ is the unit tangent vector of $\sigma$ at $x\in\sigma$.
Proposition 4.1.
[4, Lemma 12] Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact manifold $M$. There is a small constant $\varepsilon_{f}>0$ such that for any $\varepsilon_{f}$-bounded $C^{r}$ curve $\sigma:[0,1]\to M$ and any two integers $a,b\in\mathbb{Z}$, there is a family $\Theta$ of affine maps from $[0,1]$ to itself with the following properties:
  1. 1.
    $\left\{x\in\sigma:~\left\lceil\log\left\|Df_{x}\right\|\right\rceil=a,\,\left\lceil\log\left\|Df\left(\hat{x}\right)\right\|\right\rceil=b\right\}\subset\cup_{\theta\in\Theta}\sigma\circ\theta$,
  2. 2.
    $f\circ\sigma\circ\theta$ is bounded for any $\theta\in\Theta$,
  3. 3.
    $|\theta^{\prime}|\leq e^{-\frac{a-b+3r}{r-1}}$ for any $\theta\in\Theta$,
  4. 4.
    $\log\#\Theta\leq A_{r}+\frac{a-b}{r-1}$ where $A_{r}$ is a universal constant depending only on $r$.
Remark 4.2.
Compared with [4, Lemma 12], we slightly modify the third property by adding the term ” $3r$ ” in the exponent. This modification is acceptable, since the contraction factor is only required to depend on $r$, and any adjustment by a uniform constant depending solely on $r$ does not affect the validity of the result. The role of the additional term ” $3r$ ” is purely technical (see the remarks following Proposition 4.8 for details).
The following result, due to Burguet [5, Lemma 8], can be regarded as an $n$-step and local version of the reparametrization result stated in Proposition 4.1.
For $x\in M$, $n\in\mathbb{N}$ and $\varepsilon>0$, recall the usual $(n,\varepsilon)$ dynamical ball at $x$ is defined by
$B(x,n,\varepsilon):=\left\{y\in M:\ d\left(f^{k}(y),f^{k}(x)\right)<\varepsilon\ \text{ for all }0\leq k<n\right\}.$
Proposition 4.3.
[5, Lemma 8] Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact manifold $M$. For any integer $p\geq 1$, there are a small constant $\varepsilon_{f^{p}}>0$ and a large constant $B_{p}$ such that for any $\varepsilon_{f^{p}}$-bounded $C^{r}$ curve $\sigma:[0,1]\to M$, any $x\in\sigma$ and any $n\in\mathbb{N}$, there is a family $\Theta$ of affine maps from $[0,1]$ to itself with the following properties:
  1. 1.
    $B\left(x,n,\varepsilon_{f^{p}}\right)\cap\sigma\subset\cup_{\theta\in\Theta}\sigma\circ\theta$,
  2. 2.
    $f^{k}\circ\sigma\circ\theta$ is $\varepsilon_{f^{p}}$-bounded for any $\theta\in\Theta$ and $0\leq k\leq n$,
  3. 3.
[…]
    where $A_{r}$ is a universal constant depending only on $r$.
Remark 4.4.
Compared with [5, Lemma 8], we insert ceiling operations in the third property. This modification is acceptable, since the resulting discrepancy is at most linear in $n$ and can therefore be absorbed into the constant $A_{r}$. Also, to simplify notation, we assume that the constant $A_{r}$ appearing here is the same as the one in Proposition 4.1.
4.2 Decomposition of the volume growth
Recall that $\rho:\hat{M}\to\mathbb{R}$ is the continuous function
$\rho(x,[v]):=\log\left\|Df_{x}\left(v\right)\right\|$
where $v$ is a unit vector in $T_{x}M$. Define another continuous function $\rho^{\prime}:\hat{M}\to\mathbb{R}$ (introduced by Burguet [4]) by
$\rho^{\prime}(x,[v]):=\rho(x,[v])-\frac{1}{r}\log\left\|Df_{x}\right\|.$
Definition 4.5.
We call $n\geq 1$ a geometric time of $\hat{x}=(x,[v])$ if for any $0\leq k\leq n-1$,
$\sum_{i=k}^{n-1}\rho^{\prime}\left(\hat{f}^{i}(\hat{x})\right)\geq n-k.$ (11)
We let E(x^)E($\hat${x}) denote the set of all geometric times of $\hat{x}$.
[…]
Lemma 4.6.
Let $\hat{x}\in\hat{M}$. Suppose that $m\geq 1$ is not a geometric time of $\hat{x}$, and let $n<m$ be the largest geometric time smaller than $m$ (set $n=0$ if there is no geometric time before $m$). Then
$\sum_{i=n}^{m-1}\rho^{\prime}\left(\hat{f}^{i}(\hat{x})\right)<m-n.$
Proof.
Since $m$ is not a geometric time, there exists $k\in\{0,\dots,m-1\}$ such that
$\sum_{i=k}^{m-1}\rho^{\prime}\left(\hat{f}^{i}(\hat{x})\right)<m-k.$ (12)
Choose $k$ minimal with this property. If $k>n$, since $k$ is not a geometric time, there exists $\ell\in\{0,\dots,k-1\}$ such that
$\sum_{i=\ell}^{k-1}\rho^{\prime}\left(\hat{f}^{i}(\hat{x})\right)<k-\ell.$
Adding this to inequality (12) yields with $\ell<k$
$\sum_{i=\ell}^{m-1}\rho^{\prime}\left(\hat{f}^{i}(\hat{x})\right)<m-\ell,$
contradicting the minimality of $k$. Therefore we must have $k\leq n$. If $k=n$, the conclusion holds by inequality (12) (this remains true even in the case $n=0$, where $k$ must also be 0). If $k<n$, since $n$ is a geometric time,
$\sum_{i=k}^{n-1}\rho^{\prime}\left(\hat{f}^{i}(\hat{x})\right)\geq n-k.$
Subtracting this from the inequality (12) yields
$\sum_{i=n}^{m-1}\rho^{\prime}\left(\hat{f}^{i}(\hat{x})\right)<(m-k)-(n-k)=m-n,$
which is exactly the desired conclusion.
Given a subset $E\subset\mathbb{N}$ and a number $L\in\mathbb{N}$, we define
$E^{L}:=\bigcup[i,j]$
[…]
Let $\varepsilon_{f}\gg\varepsilon_{f^{p}}$ be the small constants in Proposition 4.1 w.r.t. $f$ and $f^{p}$.
Definition 4.7.
A periodic point $z$ of $f$ is called a periodic source if both Lyapunov exponents of $z$ are positive.
Let $T_{z}$ denote the period of $z$ and let $a:=\min_{i\neq j}\,d\left(f^{i}(z),f^{j}(z)\right)$. We fix a small number $0<r_{z}<\min\{\frac{a}{100},\frac{\varepsilon_{f}}{100\left\|Df\right\|}\}$  [^18] The purpose of introducing the small number $r_{z}$ will be revealed in Proposition 4.10. such that for any $0\leq k\leq T_{z}-1$,
  1. 1.
    $r_{z}=r_{f^{k}\left(z\right)}$,
  2. 2.
    $f^{-T_{z}}\left(B\left(f^{k}(z),r_{z}\right)\right)\subset B\left(f^{k}(z),r_{z}\right),$
  3. 3.
    $f\left(B\left(f^{k}(z),r_{z}\right)\right)\subset B\left(f^{k+1}(z),\frac{a}{100}\right).$
We list all periodic sources by $z_{1},z_{2},z_{3},\cdots$ (since there are at most countably many of them). We may choose $r_{z}$ sufficiently small so that the associated balls $B(z_{i},r_{z_{i}})$ are pairwise disjoint (since periodic sources form a discrete set).
Given a $C^{r}$ curve $\sigma:[0,1]\to M$ and a point $x\in\sigma$, recall that we denote by $\hat{x}=(x,[v])\in\hat{M}$ where $v$ is the unit tangent vector of $\sigma$ at $x$. Given any $L,m,n,p\in\mathbb{N}$ with $m\leq n$, any $E\subset[0,n)$ and a family of integers $\beta:=\left\{\left(\beta^{\prime}_{i},\,\beta^{\prime\prime}_{i}\right)\right\}_{i\in[0,n)}$, we define
$O^{p,m,L}_{n,E,\beta}=\Big\{x\in\sigma:$ $\left(E(\hat{x})\cap[0,n)\right)^{L}=E,\,\,\exists 1\leq s\leq n,\text{ s.t. }x\in B(z_{s},m,r_{z_{s}})\text{ and }$
$\frac{1}{p}\left\lceil\log\left\|Df^{p}_{f^{i}(x)}\right\|\right\rceil=\beta^{\prime}_{i},\,\frac{1}{p}\left\lceil\log\left\|Df^{p}\left(\hat{f}^{i}\left(\hat{x}\right)\right)\right\|\right\rceil=\beta^{\prime\prime}_{i},\,i\in[m,n)\cap E,$
$\left\lceil\log\left\|Df_{f^{i}(x)}\right\|\right\rceil=\beta^{\prime}_{i},\,\left\lceil\log\left\|Df\left(\hat{f}^{i}\left(\hat{x}\right)\right)\right\|\right\rceil=\beta^{\prime\prime}_{i},\,i\in[0,n)\setminus\left([m,n)\cap E\right)\Big\}.$
If $m=0$, we define
$O^{p,0,L}_{n,E,\beta}=\Big\{x\in\sigma:$ $\left(E(\hat{x})\cap[0,n)\right)^{L}=E,\,\,\forall 1\leq s\leq n,\text{ s.t. }x\notin B(z_{s},r_{z_{s}})\text{ and }$
$\frac{1}{p}\left\lceil\log\left\|Df^{p}_{f^{i}(x)}\right\|\right\rceil=\beta^{\prime}_{i},\,\frac{1}{p}\left\lceil\log\left\|Df^{p}\left(\hat{f}^{i}\left(\hat{x}\right)\right)\right\|\right\rceil=\beta^{\prime\prime}_{i},\,i\in[m,n)\cap E,$
$\left\lceil\log\left\|Df_{f^{i}(x)}\right\|\right\rceil=\beta^{\prime}_{i},\,\left\lceil\log\left\|Df\left(\hat{f}^{i}\left(\hat{x}\right)\right)\right\|\right\rceil=\beta^{\prime\prime}_{i},\,i\in[0,n)\setminus\left([m,n)\cap E\right)\Big\}.$
Proposition 4.8.
Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact surface $M$. Consider the following setting:
  • any $L,m,n,p\in\mathbb{N}$ with $m\leq n$,
  • any $\varepsilon_{f}$-bounded $C^{r}$ curve $\sigma:[0,1]\to M$,
  • any $O^{p,m,L}_{n,E,\beta}$ with $\beta:=\left\{\left(\beta^{\prime}_{i},\beta^{\prime\prime}_{i}\right)\right\}_{i\in[0,n)}$.
Then we have
$\log{\rm Vol}\left(f^{n}\left(O^{p,m,L}_{n,E,\beta}\right)\right)$ $\leq\#\left([m,n)\cap E\right)\cdot h_{\rm top}(f)+\left(m+\#\left([m,n)\setminus E\right)\right)\cdot\frac{\log\left\|Df\right\|}{r}$
$+\sum_{k\in[m,n)\cap E}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}+C_{n,L}.$
where the error term $C_{n,L}$ satisfies
$\limsup_{L\to+\infty}\limsup_{n\to+\infty}\frac{C_{n,L}}{n}=C_{r}$
for some universal constant $C_{r}$ depending only on $r$.
The proof of Proposition 4.8 is rather involved. We therefore first explain the underlying idea.
The set $O^{p,m,L}_{n,E,\beta}$ represents a portion of the curve $\sigma$ on which all points share the same collection of (L-extended) geometric times $E$ (we call the complementary times of $E$ the neutral times) and the same derivative data along their orbits. For simplicity, we assume that $\sigma$ intersects only one periodic source, namely a repeller $z$.
We decompose the time interval $[0,n)$ into three parts  [^19] In the proof of Proposition 4.8, an additional part, referred to as tail times, is introduced for technical reasons. : trapping times, (partial) geometric times, and (partial) neutral times,
$0,m),\quad[m,n)\cap E,\quad[m,n)\setminus E.$
  • Trapping times $[0,m)$. During this period, the set $O^{p,m,L}_{n,E,\beta}$ remains confined in a small neighborhood of the repeller $z$. Consequently, its growth is purely local  [^20] Positivity of the local volume growth is not excluded even near a repeller; it may occur when oscillatory effects dominate the geometric repelling of the region. See Example 6.8.. The corresponding growth rate is bounded by the Yomdin term $\frac{\lambda^{+}(f)}{r}$.
  • Geometric times $m,n)\cap E$. At these times, $O^{p,m,L}_{n,E,\beta}$ exhibits both a global growth, bounded by $\left([m,n)\cap E\right)\cdot h_{\rm top}(f)$, and a local growth, bounded by
    $\sum_{k\in[m,n)\cap E}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}.$
    In Proposition 4.8, we retain the latter term temporarily and later show, in Proposition 4.10, that its contribution is negligible.
  • Neutral times $m,n)\setminus E$. At neutral times, the growth of $O^{p,m,L}_{n,E,\beta}$ is again purely local, with growth rate bounded by the Yomdin term $\frac{\lambda^{+}(f)}{r}$.
It is worth emphasizing that, at neutral times, the curve does not necessarily fail to grow. Rather, by definition of geometric times, the growth rate is modest-specifically, it is bounded by a constant strictly less than $1$ (see the definition of geometric times (11)). The precise threshold is irrelevant for our purposes; what matters is that this bound is uniform and does not depend on the dynamics of $f$.
Proof of Proposition 4.8.
We divide the proof into three steps.
Step 1: An inductive construction of reparametrizations.
We first define some numbers:
  • Let $n_{E}$ be the maximal geometric time in $[0,n)$, that is, the largest element of $E\cap[0,n)$. If no such geometric time exists, we set $n_{E}=0$.
  • Let $\mathcal{A}$ be the subset of $0,n_{E}]$ obtained by removing all the integers $k$ such that
    $k,\,k+1\in[m,n_{E}]\cap E.$
    We enumerate the elements of $\mathcal{A}$ as
    $0=a_{1}<a_{2}<\cdots<a_{l_{0}}=n_{E}.$
[…]
    • Trapping times: $\Gamma^{1}_{i}:=\mathcal{A}_{i}\cap[0,m)$,
    • Geometric times  [^22] We emphasize that the set $\Gamma^{2}_{i}$ is not meant to represent all geometric times. For technical reasons, it only records the right endpoints of the geometric intervals in $m,n_{E}]\cap E$.: $\Gamma^{2}_{i}:=\left(\mathcal{A}_{i}\cap E\right)\setminus[0,m)$,
    • Neutral times: $\Gamma^{3}_{i}=\mathcal{A}_{i}\setminus\left(\Gamma^{1}_{i}\cup\Gamma^{2}_{i}\right)$,
    • Tail times: $n_{E},n)$.
Recall that $1\gg\varepsilon_{f}\gg\varepsilon_{f^{p}}>0$ and $A_{r}$ are the constants in Proposition 4.1 w.r.t. $f$ and $f^{p}$. Let $B_{p}$ be the constant in Proposition 4.3. We may further assume that $\varepsilon_{f^{p}}$ is sufficiently small and $B_{p}$ is sufficiently large such that
  • for any $k\in\mathbb{N}$, the minimal cardinality of a $(k,\varepsilon)$-spanning set of $M$ is bounded above:
    $\log r(f,k,\varepsilon_{f^{p}})\leq B_{p}+k\left(h_{\rm top}(f)+A_{r}\right),$
  • there is a family $\hat{\Theta}$  [^24] Such a family $\hat{\Theta}$ can be obtained by adapting the argument in Lemma 6.7, yielding the bound $\hat{\Theta}\lesssim\frac{\varepsilon_{f}}{\varepsilon_{f^{p}}}$. of affine maps from $[0,1]$ into itself with
    $\hat{\Theta}\leq B_{p},\quad\bigcup_{\theta\in\hat{\Theta}}\theta\left([0,1]\right)=[0,1]$
    such that for any $\varepsilon_{f}$-bounded curve $\widetilde{\sigma}$ and any $\theta\in\hat{\Theta}$, $\widetilde{\sigma}\circ\theta$ is an $\varepsilon_{f^{p}}$-bounded curve.
Define
$\Delta:=\left\{a_{k}:~a_{k+1}\in\Gamma^{2}_{l_{0}}\right\}.$
Given $1\leq s\leq n$ and a family of integers $\gamma:=\left\{\left(\gamma^{\prime}_{i},\,\gamma^{\prime\prime}_{i}\right)\right\}_{i\in[0,n)}$, we define  [^25] In the case $m=0$, we define $\Omega^{p,0,L,s}_{n,E,\gamma}$ by replacing the condition $x\in B(z_{s},m,r_{z_{s}})$” in $\Omega^{p,m,L,s}_{n,E,\gamma}$ with $x\notin B(z_{s},r_{z_{s}})$”, as before. :
$\Omega^{p,m,L,s}_{n,E,\gamma}$$\Big\{$ $x\in\sigma:~\left(E(\hat{x})\cap[0,n)\right)^{L}=E,\,x\in B(z_{s},m,r_{z_{s}})\text{ and }$
$\frac{1}{p}\left\lceil\log\left\|Df^{p}_{f^{i}(x)}\right\|\right\rceil=\gamma^{\prime}_{i},\,\frac{1}{p}\left\lceil\log\left\|Df^{p}\left(\hat{f}^{i}\left(\hat{x}\right)\right)\right\|\right\rceil=\gamma^{\prime\prime}_{i},\,i\in\left(\left([m,n)\cap E\right)\setminus\Gamma^{2}_{l_{0}}\right)\cup\Delta,$
$\left\lceil\log\left\|Df_{f^{i}(x)}\right\|\right\rceil=\gamma^{\prime}_{i},\,\left\lceil\log\left\|Df\left(\hat{f}^{i}\left(\hat{x}\right)\right)\right\|\right\rceil=\gamma^{\prime\prime}_{i},\,i\in\left(\left([0,n)\setminus\left([m,n)\cap E\right)\right)\cup\Gamma^{2}_{l_{0}}\right)\setminus\Delta\Big\}.$
The definition of $\Omega^{p,m,L,s}_{n,E,\gamma}$ differs from that of $O^{p,m,L}_{n,E,\beta}$ in two respects: first, it fixes a periodic source $z_{s}$; second, it modifies the definition at times belonging to $\Delta$ and $\Gamma^{2}_{l_{0}}$ for technical convenience. These times constitute only a small portion of $[0,n)$ (at most $2n/L$) and therefore do not affect the asymptotic estimates.
Next we carry out an inductive construction of a family of reparametrizations to cover $\Omega^{p,m,L,s}_{n,E,\gamma}$.
Claim.
There are families $\{\Theta\}_{1\leq i\leq l_{0}}$ of affine maps from $[0,1]$ to itself such that for any $1\leq i\leq l_{0}$,
  1. 1.
    $\Omega^{p,m,L,s}_{n,E,\gamma}\subset\cup_{\theta\in\Theta_{i}}\sigma\circ\theta$,
  2. 2.
    for each $\theta_{i}\in\Theta_{i}$ ($i>1$) with $a_{i}\notin\Gamma^{2}_{i}$, there are some $\theta_{i-1}\in\Theta_{i-1}$ and some affine map $\theta:[0,1]\to[0,1]$ with $|\theta^{\prime}|\leq e^{-\frac{\gamma^{\prime}_{a_{i-1}}-\gamma^{\prime\prime}_{a_{i-1}}+3r}{r-1}}$ such that $\theta_{i}=\theta_{i-1}\circ\theta$,
  3. 3.
    for any $\theta\in\Theta_{i}$, $f^{a_{i}}\circ\sigma\circ\theta$ is $\varepsilon_{f}$-bounded,
  4. 4.
    we have  [^26] By convention, we set $a,b)=\emptyset$ if $a>b$.
    $\log\#\Theta_{i}\leq\#\left([m,a_{i}]\cap E\right)\cdot h_{\rm top}(f)+\sum_{k<a_{i}}\frac{\gamma^{\prime}_{k}-\gamma^{\prime\prime}_{k}}{r-1}+\#\Gamma^{2}_{i}\cdot 3B_{p}+a_{i}\cdot 2A_{r}$
    where $A_{r}$ is the constant given in Proposition 4.1, which coincides with that in Proposition 4.3.
Proof of Claim.
We start an inductive construction to build $\Theta_{0\leq i\leq l_{0}}$.
Initial step:
For $i=1$ ($a_{1}=0$), the curve $\sigma$ is $\varepsilon_{f}$-bounded by our assumption. We define $\Theta_{1}:=\{{\rm Id:[0,1]\to[0,1]}\}$. The properties in the claim are trivially satisfied.
Inductive step:
Assume that we have families $\{\Theta_{k}\}_{1\leq k\leq i}$ of affine maps satisfying the four properties in the claim. We consider three cases:
  1. 1.
    $a_{i+1}\in\Gamma^{1}_{i+1}$.
    Note that in this case,
    $a_{i+1}=a_{i}+1=i\in[0,m).$
    As a consequence, given any $\theta_{i}\in\Theta_{i}$, after composing with a suitable reparametrization, we may assume  [^27] This follows directly from the definition of $\Omega^{p,m,L,s}_{n,E,\gamma}$, since we only consider points lying in the dynamical ball $B(z_{s},m,r_{z_{s}})$ where $r_{z_{s}}<\frac{\varepsilon_{f}}{100\left\|Df\right\|}$.
    $f^{a_{i}}\circ\sigma\circ\theta_{i}\subset B\left(f^{a_{i}}\left(z_{s}\right),\frac{\varepsilon_{f}}{10\left\|Df\right\|}\right).$ (13)
    Applying Proposition 4.1 with $a=\gamma^{\prime}_{a_{i}},b=\gamma^{\prime\prime}_{a_{i}}$, we get a family $\Theta$ of affine maps from $[0,1]$ to itself with
    $\log\#\Theta\leq\frac{\gamma^{\prime}_{a_{i}}-\gamma^{\prime\prime}_{a_{i}}}{r-1}+A_{r}$
    such that each $f^{a_{i+1}}\circ\sigma\circ\theta_{i}\circ\theta,\,\theta\in\Theta$ is bounded and
    $|\theta^{\prime}|\leq e^{-\frac{\gamma^{\prime}_{a_{i}}-\gamma^{\prime\prime}_{a_{i}}+3r}{r-1}}.$
    By (13),
    $f^{a_{i+1}}\circ\sigma\circ\theta_{i}\circ\theta\subset B\left(f^{a_{i+1}}\left(z_{s}\right),\frac{\varepsilon_{f}}{10}\right).$
    Hence we may assume, after composing $\theta$ with a suitable reparametrization  [^28] This can be shown directly from the definitions; see [4, Lemma 13] and the references therein for a precise argument., that $f^{a_{i+1}}\circ\sigma\circ\theta_{i}\circ\theta$ is $\varepsilon_{f}$-bounded. We then define
    $\Theta_{i+1}:=\{\theta_{i}\circ\theta:~\,\theta\in\Theta,\,\theta_{i}\in\Theta_{i}\}.$
    Note in this case,
    $\Gamma^{1}_{i+1}=\Gamma^{1}_{i}\cup\{a_{i+1}\},\quad\Gamma^{2}_{i+1}=\Gamma^{2}_{i},\quad\Gamma^{3}_{i+1}=\Gamma^{3}_{i},\quad a_{i+1}-a_{i}=1,$
    and
    $\left([m,a_{i}]\cap E\right)=\#\left([m,a_{i+1}]\cap E\right)=0.$
    Hence we have the desired bound on $\log\#\Theta_{i+1}$:
    $\log\#\Theta_{i+1}$ $\leq\log\#\Theta_{i}+\#\log\Theta$
    $\leq\#\left([m,a_{i+1}]\cap E\right)\cdot h_{\rm top}(f)+\sum_{k<a_{i+1}}\frac{\gamma^{\prime}_{k}-\gamma^{\prime\prime}_{k}}{r-1}+\#\Gamma^{2}_{i+1}\cdot 3B_{p}+a_{i+1}\cdot 2A_{r}.$
    The remaining properties asserted in the claim are obvious.
  2. 2.
    $a_{i+1}\in\Gamma^{2}_{i+1}$.
    In this case, $a_{i}\notin\Gamma^{2}_{i}$ and $(a_{i},a_{i+1}]\subset E$.
    Recall the definitions of $\varepsilon_{f^{p}}$, $B_{p}$ and $\hat{\Theta}$ ($\hat{\Theta}\leq B_{p}$) introduced before the claim.
    For any $\theta_{i}\in\Theta_{i}$ and $\hat{\theta}\in\hat{\Theta}$, $f^{a_{i}}\circ\sigma\circ\theta_{i}\circ\hat{\theta}$ is an $\varepsilon_{f^{p}}$-bounded curve. Next, we decompose each $f^{a_{i}}\circ\sigma\circ\theta_{i}\circ\hat{\theta}$ into smaller pieces, with the decomposition taken up to time $a_{i+1}-a_{i}$. We choose some $(a_{i+1}-a_{i},\varepsilon_{f^{p}})$ spanning set $\Omega\subset f^{a_{i}}\circ\sigma\circ\theta_{i}\circ\hat{\theta}$ with
    $\log\#\Omega=\log r(f,a_{i+1}-a_{i},\varepsilon_{f^{p}})\leq B_{p}+\left(a_{i+1}-a_{i}\right)\cdot\left(h_{\rm top}(f)+A_{r}\right).$
    Note that we are only concerned with points $z\in f^{a_{i}}\circ\sigma\circ\theta_{i}\circ\hat{\theta}$ satisfying
    $\frac{1}{p}\left\lceil\log\left\|Df^{p}_{f^{j}(z)}\right\|\right\rceil=\gamma^{\prime}_{a_{i}+j},\quad\frac{1}{p}\left\lceil\log\left\|Df^{p}\left(\hat{f}^{i}\left(\hat{z}\right)\right)\right\|\right\rceil=\gamma^{\prime\prime}_{a_{i}+j},\quad j\in[0,a_{i+1}-a_{i}).$
    where $\hat{z}:=(z,[v])$, with $v$ denoting the unit tangent vector to the curve $f^{a_{i}}\circ\sigma\circ\theta_{i}\circ\hat{\theta}$ at $z$. Hence, without loss of generality, we many assume that all points in $\Omega$ satisfy these relations.
    Applying Proposition 4.3 to each point $z\in\Omega$ with $n=a_{i+1}-a_{i}$, we have a family $\Theta_{z}$ of affine maps from $[0,1]$ to itself with the following properties:
    1. (a)
      $B(z,a_{i+1}-a_{i},\varepsilon_{f^{p}})\cap f^{a_{i}}\circ\sigma\circ\theta_{i}\circ\hat{\theta}\subset\cup_{\theta\in\Theta_{z}}f^{a_{i}}\circ\sigma\circ\theta_{i}\circ\hat{\theta}\circ\theta$,
    2. (b)
      $f^{k}\circ f^{a_{i}}\circ\sigma\circ\theta_{i}\circ\hat{\theta}\circ\theta$ is $\varepsilon_{f^{p}}$-bounded for any $\theta\in\Theta_{z}$ and $0\leq k\leq a_{i+1}-a_{i}$,
    3. (c)
      We have
      $\log\#\Theta_{z}\leq B_{p}+\left(a_{i+1}-a_{i}\right)\cdot A_{r}+\frac{1}{r-1}\cdot\sum_{k=a_{i}}^{a_{i+1}-1}\left(\gamma^{\prime}_{k}-\gamma^{\prime\prime}_{k}\right).$
    Note that
    $\Gamma^{1}_{i+1}=\Gamma^{1}_{i},\quad\Gamma^{2}_{i+1}=\Gamma^{2}_{i}\cup\{a_{i+1}\},\quad\Gamma^{3}_{i+1}=\Gamma^{3}_{i}.$
    Since the half-open interval $(a_{i},a_{i+1}]\subset E$, we have
    $\left([m,a_{i+1}]\cap E\right)=\#\left([m,a_{i}]\cap E\right)+a_{i+1}-a_{i}.$
    We then define
    $\Theta_{i+1}:=\left\{\theta_{i}\circ\hat{\theta}\circ\theta:~\theta_{i}\in\Theta_{i},\,\hat{\theta}\in\hat{\Theta},\,\theta\in\Theta_{z},\,z\in\Omega\right\}$
    and we have
    $\log\#\Theta_{i+1}$ $\leq\log\#\Theta_{i}+\log\#\hat{\Theta}+\max_{z\in\Omega}\log\#\Theta_{z}+\log\#\Omega$
    $\leq\#\left([m,a_{i+1})\cap E\right)\cdot h_{\rm top}(f)+\sum_{k<a_{i+1}}\frac{\gamma^{\prime}_{k}-\gamma^{\prime\prime}_{k}}{r-1}+\#\Gamma^{2}_{i+1}\cdot 3B_{p}+a_{i+1}\cdot 2A_{r}.$
    The remaining properties asserted in the claim are obvious. The terms $3B_{p},2A_{r}$” arise from this case.
  3. 3.
    $a_{i+1}\in\Gamma^{3}_{i+1}$.
    In this case, $a_{i+1}-a_{i}=1$.
    Given any $\theta_{i}\in\Theta_{i}$, for the $\varepsilon_{f}$-bounded curve $f^{a_{i}}\circ\sigma\circ\theta_{i}$, we apply Proposition 4.1 with $a=\gamma^{\prime}_{a_{i}},b=\gamma^{\prime\prime}_{a_{i}}$ and we get a family $\Theta$ of affine maps from $[0,1]$ to itself with
    $\log\#\Theta\leq\frac{\gamma^{\prime}_{a_{i}}-\gamma^{\prime\prime}_{a_{i}}}{r-1}+A_{r}$
    such that each $f^{a_{i+1}}\circ\sigma\circ\theta_{i}\circ\theta,\,\theta\in\Theta$ is bounded and
    $|\theta^{\prime}|\leq e^{-\frac{\gamma^{\prime}_{a_{i}}-\gamma^{\prime\prime}_{a_{i}}+3r}{r-1}}.$
    Next we show that $f^{a_{i+1}}\circ\sigma\circ\theta_{i}\circ\theta,\,\theta\in\Theta$ is actually $\varepsilon_{f}$-bounded. Define
    $j=\max\left\{k:~a_{k}\in E,\,\,a_{k}<a_{i+1}\right\}.$
    If the set in the right side is empty, we set $j=1$ which gives $a_{j}=a_{1}=0$. We also note that $(a_{j},a_{i+1}]\cap E=\emptyset$ by the maximality of $j$. Consequently,
    $a_{i+1}-a_{j}=i+1-j.$
[…]
    $\sum_{k=a_{j}}^{a_{i+1}-1}\rho^{\prime}\left(\hat{f}^{k}(\hat{x})\right)=\sum_{k=a_{j}}^{a_{i}}\rho^{\prime}\left(\hat{f}^{k}(\hat{x})\right)=\sum_{k=j}^{i}\rho^{\prime}\left(\hat{f}^{a_{k}}(\hat{x})\right)<a_{i+1}-a_{j}=i+1-j.$
    Take the ceiling operation into consideration, we then have
    $\sum_{k=j}^{i}\gamma^{\prime\prime}_{a_{k}}-\frac{\gamma^{\prime}_{a_{k}}}{r}<2\left(a_{i+1}-a_{j}\right)=2\left(i+1-j\right).$ (14)
    Let $t\in[0,1]$ be such that $\sigma\circ\theta_{i}\circ\theta(t)=x\in\Omega^{p,m,L,s}_{n,E,\gamma}$ for some $x$. By the second assumption in the claim, there are some $\theta_{j}\in\Theta_{j}$ and some affine maps $\widetilde{\theta}_{j},\widetilde{\theta}_{j+1},\cdots,\widetilde{\theta}_{i-1}$ with
    $|\widetilde{\theta}^{\prime}_{k}|\leq e^{-\frac{\gamma^{\prime}_{a_{k}}-\gamma^{\prime\prime}_{a_{k}}+3r}{r-1}},\quad j\leq k\leq i-1$
    such that
    $\theta_{i}=\theta_{j}\circ\widetilde{\theta}_{j}\circ\widetilde{\theta}_{j+1}\circ\cdots\circ\widetilde{\theta}_{i-1}.$
    We have
[…]
    $\leq 2\left\|Df^{a_{i+1}-a_{j}}\left(\hat{f}^{a_{j}}(\hat{x})\right)\right\|\cdot\left\|d_{t}\left(f^{a_{j}}\circ\sigma\circ\theta_{i}\circ\theta\right)\right\|$
    $\leq 2\left\|Df^{a_{i+1}-a_{j}}\left(\hat{f}^{a_{j}}(\hat{x})\right)\right\|\cdot\left\|d\left(f^{a_{j}}\circ\sigma\circ\theta_{i}\right)\right\|\cdot|\theta^{\prime}|$
    $\leq 2\left\|Df^{a_{i+1}-a_{j}}\left(\hat{f}^{a_{j}}(\hat{x})\right)\right\|\cdot\left\|d\left(f^{a_{j}}\circ\sigma\circ\theta_{j}\right)\right\|\cdot\left(\prod_{k=j}^{i-1}|\widetilde{\theta}^{\prime}_{k}|\right)\cdot|\theta^{\prime}|$
    $\leq 2\exp{\left(\sum_{k=j}^{i}\gamma^{\prime\prime}_{a_{k}}\right)}\cdot\varepsilon_{f}\cdot\exp{\left(\sum_{k=j}^{i}-\frac{\gamma^{\prime}_{a_{k}}-\gamma^{\prime\prime}_{a_{k}}+3r}{r-1}\right)}$
    $=2\varepsilon_{f}\cdot\exp{\left(\sum_{k=j}^{i}\gamma^{\prime\prime}_{a_{k}}-\frac{\gamma^{\prime}_{a_{k}}-\gamma^{\prime\prime}_{a_{k}}+3r}{r-1}\right)}$
    $=2\varepsilon_{f}\cdot\exp{\left(\sum_{k=j}^{i}\frac{r\cdot\left(\gamma^{\prime\prime}_{a_{k}}-\frac{\gamma^{\prime}_{a_{k}}}{r}\right)-3r}{r-1}\right)}$
    $=2\varepsilon_{f}\cdot\exp{\left(\frac{r}{r-1}\cdot\left(\sum_{k=j}^{i}\gamma^{\prime\prime}_{a_{k}}-\frac{\gamma^{\prime}_{a_{k}}}{r}\right)-\frac{3r\left(i-j+1\right)}{r-1}\right)}$
[…]
    $\leq 2\varepsilon_{f}\cdot e^{-1}$
    $\leq\varepsilon_{f}.$
    This shows that $f^{a_{i+1}}\circ\sigma\circ\theta_{i}\circ\theta$ is not only bounded, but $\varepsilon_{f}$-bounded. One can also see from the above estimates the technical role played by the term ” $3r$ ” in Proposition 4.1.
    We define $\Theta_{i+1}:=\{\theta_{i}\circ\theta:~\theta_{i}\in\Theta_{i},\theta\in\Theta\}$. Then we have
    $\Gamma^{1}_{i+1}=\Gamma^{1}_{i},\quad\Gamma^{2}_{i+1}=\Gamma^{2}_{i},\quad\Gamma^{3}_{i+1}=\Gamma^{3}_{i}\cup\{a_{i+1}\},\quad a_{i+1}-a_{i}=1.$
    Note
    $\left([m,a_{i+1}]\cap E\right)=\#\left([m,a_{i}]\cap E\right).$
    Hence we have the desired bound on $\log\#\Theta_{i+1}$:
    $\log\#\Theta_{i+1}$ $\leq\log\#\Theta_{i}+\#\log\Theta$
    $\leq\#\left([m,a_{i+1}]\cap E\right)\cdot h_{\rm top}(f)+\sum_{k<a_{i+1}}\frac{\gamma^{\prime}_{k}-\gamma^{\prime\prime}_{k}}{r-1}+\#\Gamma^{2}_{i+1}\cdot 3B_{p}+a_{i+1}\cdot 2A_{r}.$
    The remaining properties asserted in the claim are obvious.
The proof of the claim is complete. ∎
We define a subset of $O^{p,m,L}_{n,E,\beta}$ associated with a fixed periodic source $z_{s}$ by [^30] In the case $m=0$, we define $O^{p,0,L,s}_{n,E,\gamma}$ by replacing the condition $x\in B(z_{s},m,r_{z_{s}})$” in $O^{p,m,L,s}_{n,E,\gamma}$ with $x\notin B(z_{s},r_{z_{s}})$”, as before.
$O^{p,m,L,s}_{n,E,\beta}=\Big\{x\in\sigma:$ $\left(E(\hat{x})\cap[0,n)\right)^{L}=E,\,\,x\in B(z_{s},m,r_{z_{s}})\text{ and }$
$\frac{1}{p}\left\lceil\log\left\|Df^{p}_{f^{i}(x)}\right\|\right\rceil=\beta^{\prime}_{i},\,\frac{1}{p}\left\lceil\log\left\|Df^{p}\left(\hat{f}^{i}\left(\hat{x}\right)\right)\right\|\right\rceil=\beta^{\prime\prime}_{i},\,i\in[m,n)\cap E,$
$\left\lceil\log\left\|Df_{f^{i}(x)}\right\|\right\rceil=\beta^{\prime}_{i},\,\left\lceil\log\left\|Df\left(\hat{f}^{i}\left(\hat{x}\right)\right)\right\|\right\rceil=\beta^{\prime\prime}_{i},\,i\in[0,n)\setminus\left([m,n)\cap E\right)\Big\}.$
We note that by definition,
$O^{p,m,L,s}_{n,E,\beta}\subset\bigcup_{\gamma}\Omega^{p,m,L,s}_{n,E,\gamma}$ (15)
where the union is taken over all possible $\gamma:=\left\{\left(\gamma^{\prime}_{i},\,\gamma^{\prime\prime}_{i}\right)\right\}_{i\in[0,n)}$ such that
$\gamma^{\prime}_{i}=\beta^{\prime}_{i},\quad\gamma^{\prime\prime}_{i}=\beta^{\prime\prime}_{i},\quad i\in[0,n)\setminus\left(\Delta\cup\Gamma_{l_{0}}^{2}\right).$
There are at most
$\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)^{\frac{2n}{L}}$
such $\gamma$, since both $\Delta$ and $\Gamma_{l_{0}}^{2}$ have cardinality at most $n/L$. Indeed, the set $\Gamma_{l_{0}}^{2}$ records the right endpoints of geometric intervals, and $\Delta$ has the same cardinality as $\Gamma_{l_{0}}^{2}$. The number of geometric intervals (including isolated ones) is relatively small-at most $n/L$-because the complement of $E$ in $[0,n)$ is a disjoint union of intervals whose lengths are all larger than $L$.
As a consequence of (15) and the claim above, there is a family $\widetilde{\Theta}_{l_{0}}$ of affine maps from $[0,1]$ to itself such that
  1. 1.
    $O^{p,m,L,s}_{n,E,\beta}\subset\cup_{\theta\in\widetilde{\Theta}_{l_{0}}}\sigma\circ\theta$,
  2. 2.
    for any $\theta\in\widetilde{\Theta}_{l_{0}}$, $f^{n_{E}}\circ\sigma\circ\theta$ is $\varepsilon_{f}$-bounded,
  3. 3.
    recalling $a_{l_{0}}=n_{E}$,
    $\log\#\widetilde{\Theta}_{l_{0}}$ $\leq\#\left([m,n)\cap E\right)\cdot h_{\rm top}(f)+\sum_{k\in[0,n_{E})}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}+\frac{n}{L}\cdot 3B_{p}+n\cdot 2A_{r}$ (16)
    $\quad+\frac{2n}{L}\cdot\log\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)+\frac{2n}{L}\cdot\frac{\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2}{r-1}.$
    Note that the last term above
    $\frac{2n}{L}\cdot\frac{\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2}{r-1}$
    is introduced to to compensate for the discrepancy between (at the times $\Delta\cup\Gamma_{l_{0}}^{2}$)
    $\sum_{k\in[0,n_{E})}\frac{\gamma^{\prime}_{k}-\gamma^{\prime\prime}_{k}}{r-1}\quad\text{ and}\quad\sum_{k\in[0,n_{E})}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}.$
Step 2: Estimates along the time decomposition
Next, we estimate from above the quantity:
$\sum_{k\in[0,n_{E})}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}.$
  • Geometric Bound:
    The estimation of the geometric component $\sum_{k\in[m,n)\cap E}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}$ is rather delicate. We will show in Proposition 4.10 that this term is negligible, using a measure-theoretic argument. We retain this term in Proposition 4.8.
  • Neutral Bound:
    Recall that $n_{E}$ is the maximal geometric time in $[0,n)$ (if no such geometric time exists, we set $n_{E}=0$). We consider the times in $0,n_{E})\setminus E$  [^31] This part contains the neutral component (i.e., the non-geometric times $0,m)\setminus E$) of the trapping times $[0,m)$ and is treated here for technical convenience. The estimation of the remaining, geometric component (i.e., $0,m)\cap E$) will be carried out in Trapping Bound below.. We decompose $0,n_{E})\setminus E$ into maximal disjoint intervals:
    $0,n_{E})\setminus E=\bigcup[c_{i},d_{i}).$
    We note that $d_{i}$ must be a geometric time by maximality and consequently, for any $x\in O^{p,m,L,s}_{n,E,\beta}$,
    $0<d_{i}-c_{i}\leq\sum_{k\in[c_{i},d_{i})}\rho^{\prime}\left(\hat{f}^{k}(\hat{x})\right)\leq 10(d_{i}-c_{i})+\sum_{k\in[c_{i},d_{i})}\beta^{\prime\prime}_{k}-\frac{\beta^{\prime}_{k}}{r}.$
    Here the term $10(d_{i}-c_{i})$” is introduced to compensate for the discrepancy between the ceiling operation and the actual value. The above inequalities imply
    $\sum_{k\in[c_{i},d_{i})}-\beta^{\prime\prime}_{k}\leq 10(d_{i}-c_{i})-\sum_{k\in[c_{i},d_{i})}\frac{\beta^{\prime}_{k}}{r}.$
    Adding up all the intervals, we get
    $\sum_{k\in[0,n_{E})\setminus E}-\beta^{\prime\prime}_{k}\leq 10\#\left([0,n_{E})\setminus E\right)-\sum_{k\in[0,n_{E})\setminus E}\frac{\beta^{\prime}_{k}}{r}.$
    Hence,
    $\sum_{k\in[0,n_{E})\setminus E}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}$ $\leq\frac{10\#\left([0,n_{E})\setminus E\right)}{r-1}+\sum_{k\in[0,n_{E})\setminus E}\frac{\beta^{\prime}_{k}}{r}$
    $\leq\frac{10\#\left([0,n_{E})\setminus E\right)}{r-1}+\#\left([0,n_{E})\setminus E\right)\cdot\frac{\log\left\|Df\right\|}{r}$
    $=\#\left([0,n_{E})\setminus E\right)\cdot\left(\frac{10}{r-1}+\frac{\log\left\|Df\right\|}{r}\right).$
  • Trapping Bound:
    The estimate of $\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}$ for all $k\in[0,m)\setminus E$ has already been addressed in the Neutral Bound above. It remains to treat the case where $k\in[0,m)\cap E$. We decompose $0,m)\cap E$ into maximal disjoint intervals:
    $0,m)\cap E=\left(\bigcup[c_{i},d_{i}]\right)\bigcup[m_{E},m-1]$
    where $m_{E},m-1]$ denotes the last interval if $m-1,m\in E$ (which means that the geometric interval is truncated at m1m-1). For each interval $c_{i},d_{i})$, by an almost identical argument as in the Neutral Bound, we have
    $\sum_{k\in[c_{i},d_{i})}-\beta^{\prime\prime}_{k}\leq 10(d_{i}-c_{i})-\sum_{k\in[c_{i},d_{i})}\frac{\beta^{\prime}_{k}}{r}.$
    Adding up all the intervals, we get
    $\sum_{i}\sum_{k\in[c_{i},d_{i})}-\beta^{\prime\prime}_{k}\leq 10m-\sum_{i}\sum_{k\in[c_{i},d_{i})}\frac{\beta^{\prime}_{k}}{r}.$
    For the last interval $m_{E},m-1]$, there is some geometric time $j\in(m-1-L,m-1]$. As a consequence,
    $\sum_{k\in[m_{E},m-1]}-\beta^{\prime\prime}_{k}\leq 10m-\sum_{k\in[m_{E},m-1]}\frac{\beta^{\prime}_{k}}{r}+L\cdot\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right).$ (17)
    Hence,
    $\quad\sum_{k\in[0,m)\cap E}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}$
    $=\sum_{i}\sum_{k\in[c_{i},d_{i})}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}+\sum_{d_{i}}\frac{\beta^{\prime}_{d_{i}}-\beta^{\prime\prime}_{d_{i}}}{r-1}+\sum_{k\in[m_{E},m-1]}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}$
    $\leq\frac{20m}{r-1}+\sum_{i}\sum_{k\in[c_{i},d_{i})}\frac{\beta^{\prime}_{k}}{r}+\sum_{d_{i}}\frac{\beta^{\prime}_{d_{i}}-\beta^{\prime\prime}_{d_{i}}}{r-1}+\sum_{k\in[m_{E},m-1]}\frac{\beta^{\prime}_{k}}{r}+L\cdot\frac{\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2}{r-1}$
    $\leq\#\left([0,m)\cap E\right)\cdot\frac{\log\left\|Df\right\|}{r}$
    $\quad+\frac{20m}{r-1}+\frac{m}{L}\cdot\frac{\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2}{r-1}+L\cdot\frac{\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2}{r-1}.$
    Note that, in the last step, we use the fact that there are relatively few (at most $m/L$) right endpoints $d_{i}$ among the maximal intervals in $0,m)\cap E$, since the complement of $E$ in $[0,m)$ is a disjoint union of intervals whose lengths exceed $L$ (except possibly for the first and the last ones).
  • Tail Bound:
    For the times in $n_{E},n)$, if $n_{E}<n-1$ which implies that n1n-1 is not a geometric time, for any $x\in O^{p,m,L,s}_{n,E,\beta}$, by Lemma 4.6,
    $\sum_{k\in[n_{E},n-1)}\rho^{\prime}\left(\hat{f}^{k}(\hat{x})\right)<n-n_{E}.$
    Taking the ceiling operation into consideration, we then have
    $\sum_{k\in[n_{E},n)}\beta^{\prime\prime}_{k}-\frac{\beta^{\prime}_{k}}{r}\leq 2n+10\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right).$ (18)
[…]
    For the tail times, instead of estimating $\sum_{k\in[n_{E},n)}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}$, we directly consider the volume growth. Given any reparametrization $\theta\in\widetilde{\Theta}_{l_{0}}$, recalling the last paragraph in Step 1, $f^{n_{E}}\circ\sigma\circ\theta$ is an $\varepsilon_{f}$-bounded curve and consequently, its volume (length) is less than 1. Define  [^32] We ignore the value at time $n_{E}$ in the definition of $\Omega_{\theta}$ for technical convenience, which results in a slightly larger set. Since this omission concerns only a single time instance, it does not affect the asymptotic regime.
    $\Omega_{\theta}=\Big\{t\in[0,1]:$ $x=\sigma\circ\theta(t),\quad\left\lceil\log\left\|Df_{f^{k}(x)}\right\|\right\rceil=\beta^{\prime}_{k},$
    $\left\lceil\log\left\|Df\left(\hat{f}^{k}\left(\hat{x}\right)\right)\right\|\right\rceil=\beta^{\prime\prime}_{k},\quad k\in(n_{E},n)\Big\}.$
    Note that our focus is solely on the restriction $\sigma\circ\theta|_{\Omega_{\theta}}$. By (18), we have
    $\quad\log{\rm Vol}\left(f^{n}\circ\sigma\circ\theta|_{\Omega_{\theta}}\right)$
[…]
    Note that the term $10\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)$” in the first inequality is introduced to compensate for the mismatch at time $n_{E}$.
Step 3: Conclusion.
By the estimates in Step 2,
$\sum_{k\in[0,n_{E})}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}$ $=\sum_{k\in[m,n_{E})\cap E}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}+\sum_{k\in[0,n_{E})\setminus E}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}+\sum_{k\in[0,m)\cap E}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}$
$\leq\sum_{k\in[m,n_{E})\cap E}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}+\left(m+\#\left([m,n_{E})\setminus E\right)\right)\cdot\frac{\log\left\|Df\right\|}{r}$
$\quad+\frac{30n}{r-1}+\frac{n}{L}\cdot\frac{\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2}{r-1}+L\cdot\frac{\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2}{r-1}.$
Since $\varepsilon_{f}$-bounded curve has volume (length) less than 1, by the Tail Bound in Step 2,
$\log{\rm Vol}\left(f^{n}\left(O^{p,m,L,s}_{n,E,\beta}\right)\right)\leq\log\#\widetilde{\Theta}_{l_{0}}+\left(n-n_{E}\right)\cdot\frac{\log\left\|Df\right\|}{r}+2n+20\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right).$
By the definition of $n_{E}$,
$m+\#\left([m,n_{E})\setminus E\right)+n-n_{E}=m+\#\left([m,n)\setminus E\right).$
Hence, by (16) we have
$\log{\rm Vol}\left(f^{n}\left(O^{p,m,l}_{n,E,\beta}\right)\right)$ $\leq\log n+\max_{1\leq s\leq n}\log{\rm Vol}\left(f^{n}\left(O^{p,m,L,s}_{n,E,\beta}\right)\right)$
[…]
$\leq\#\left([m,n)\cap E\right)\cdot h_{\rm top}(f)+\left(m+\#\left([m,n)\setminus E\right)\right)\cdot\frac{\log\left\|Df\right\|}{r}+\sum_{k\in[m,n)\cap E}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}$
$\quad+\frac{n}{L}\cdot 3B_{p}+n\cdot\left(2A_{r}+3\right)+\frac{30n}{r-1}$
$\quad+\left(\frac{3n}{L}\cdot\frac{1}{r-1}+\frac{L}{r-1}+\frac{2n}{L}+20\right)\cdot\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right).$
Define
$C_{n,L}:=\frac{n}{L}\cdot 3B_{p}+n\cdot\left(2A_{r}+3\right)+\frac{30n}{r-1}+\left(\frac{3n}{L}\cdot\frac{1}{r-1}+\frac{L}{r-1}+\frac{2n}{L}+20\right)\cdot\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)$
and
$C_{r}:=2A_{r}+3+\frac{30}{r-1}$
and note that
$\limsup_{L\to+\infty}\limsup_{n\to+\infty}\frac{C_{n,L}}{n}=C_{r}.$
The proof of Proposition 4.8 is then complete.
4.3 Estimation on the geometric component and proof of Theorem 2.1
Recall that given a subset $E\subset\mathbb{N}$ and a number $L\in\mathbb{N}$, we defined
$E^{L}:=\bigcup[i,j]$
where the union runs over all $i,j\in E$ with $0\leq j-i\leq L$.
Given a subset $E\subset[0,n)$ (assume that $n$ is given in advance), we write
$\delta_{E}\left(\hat{x}\right):=\frac{1}{n}\sum_{i\in E}\delta_{\hat{f}^{i}(\hat{x})}$
and we let $\delta_{E}\left(x\right)$ denote the corresponding projection, i.e.,
$\delta_{E}\left(x\right):=\frac{1}{n}\sum_{i\in E}\delta_{f^{i}(x)}.$
The following proposition shows that limits of empirical measures along geometric times are $\hat{f}-$invariant and project to invariant measures with at least one positive Lyapunov exponent.
Recall the geometric times in Definition 11 and the maximal Lyapunov exponent of $x$:
$\lambda^{+}(f,x):=\limsup_{n\to+\infty}\frac{1}{n}\log\left\|Df^{n}_{x}\right\|.$
Proposition 4.9.
Let $f$ be a $C^{1}$ diffeomorphism on a compact surface $M$. We consider a doubly indexed family of points $\hat{x}_{n,p}\in\hat{M}$ together with a family of geometric time sets $E_{n}\subset[0,n)$ (common to all indices $p$). Let $0\leq m_{n}\leq n$ be a sequence of integers and assume
$\delta_{[m_{n},n)\cap E_{n}^{L}}\left(\hat{x}_{n,p}\right)\xrightarrow{\text{$\quad n\quad$}}\hat{\mu}^{L}_{p},\quad\quad\hat{\mu}^{L}_{p}\xrightarrow{\text{$\quad L\quad$}}\hat{\mu}_{p},\quad\quad\hat{\mu}_{p}\xrightarrow{\text{$\quad p\quad$}}\hat{\mu}$
for some Borel measures (not necessarily probability measures) $\hat{\mu}^{L}_{p},\hat{\mu}_{p}$ and $\hat{\mu}$. Then $\hat{\mu}_{p}$ and $\hat{\mu}$ are $\hat{f}-$invariant. Moreover, for the projection $\mu$-almost every $x$, $\lambda^{+}(f,x)>0$ (assuming $\mu$ is not zero measure) and $\hat{\mu}$ is an unstable lift of $\mu$.
Proof.
Given $L\in\mathbb{N}$, we decompose $m_{n},n)\cap E_{n}^{L}$ into maximal disjoint intervals:
$m_{n},n)\cap E_{n}^{L}=\bigcup_{i\geq 1}[a_{i},b_{i}).$
Note that there are relatively few (at most $n/L$) such intervals $a_{i},b_{i})$, as the complement of $E^{L}_{n}$ in $[0,n)$ consists of a disjoint union of intervals with lengths exceeding $L$, except possibly the first and the last intervals. Consequently, given any continuous function $h:\hat{M}\to\mathbb{R}$ and any $n,p\in\mathbb{N}$,
[…]
Passing to the limit in $n$, we have
[…]
Then passing to the limit in $L$ gives the $\hat{f}-$invariance of $\hat{\mu}_{p}$ and the $f$-invariant of $\mu_{p}$. As a consequence, $\hat{\mu}$ and $\mu$ are also invariant.
Next we show that for $\mu$-almost every $x$, $\lambda^{+}(f,x)>0$. We first note the following two facts:
  • Since $E_{n}^{L}\subset E_{n}^{L+1}$, the above convergence w.r.t. $L$ (i.e., $\hat{\mu}^{L}_{p}\xrightarrow{\text{$\quad L\quad$}}\hat{\mu}_{p}$) is an increasing convergence where increasing means for any subset $A$,
    $\hat{\mu}^{L}_{p}(A)\leq\hat{\mu}^{L+1}_{p}(A).$
  • By the convergences of the measures, we have
[…]
    where $\alpha^{L},\alpha\in[0,1]$ are defined (independent of $p$) by
    $\alpha^{L}:=\hat{\mu}^{L}_{p}(\hat{M})=\mu^{L}_{p}(M),\quad\alpha:=\hat{\mu}_{p}(\hat{M})=\hat{\mu}(\hat{M})=\mu_{p}(M)=\mu(M).$
    Since we assume that $\mu$ is not zero measure, we may assume $\alpha^{L},\alpha>0$. Also note that $\alpha^{L}$ converges increasingly to $\alpha$.
Note that by definition, for any point $\hat{x}_{n,p}$ and any integer $i$ with $i,i+1\in E_{n}^{L}$  [^33] This additional condition $i+1\in E_{n}^{L}$ is a purely technical adjustment. It compensates for a minor asymmetry in the definition of geometric times: although $n$ itself may be declared a geometric time, the defining inequality only involves sums up to time n1n-1 (see Definition 11). Requiring $i+1\in E_{n}^{L}$ avoids boundary issues caused by this convention and does not affect any asymptotic estimates., there is $1\leq l\leq L$ such that $l+i$ is a geometric time of $\hat{x}$. We then have (recalling $\rho(x,[v]):=\log\left\|Df_{x}\left(v\right)\right\|$)
$\hat{f}^{i}(\hat{x}_{n,p})\in\Omega^{L}:=\left\{\hat{x}\in\hat{M}:~\exists 1\leq l\leq L\text{ such that }\sum_{m=0}^{l-1}\rho\left(\hat{f}^{m}\left(\hat{x}\right)\right)\geq l\right\}.$
Note that in the definition of $\Omega^{L}$ above, we use the larger function $\rho$ rather than $\rho^{\prime}$ for simplicity (since $\rho\geq\rho^{\prime}$). Note that the number of indices $i$ satisfying $i\in E_{n}^{L}$ and $i+1\notin E_{n}^{L}$ is at most $\frac{n}{L}$, since each such $i$ corresponds to the right endpoint of a maximal geometric interval and the complement of $E_{n}^{L}$ consists of intervals of length at least $L$. As a consequence, for any $n,p\in\mathbb{N}$, we have
$\big|\delta_{[m_{n},n)\cap E_{n}^{L}}\left(\hat{x}_{n,p}\right)\left(\Omega^{L}\right)-\delta_{[m_{n},n)\cap E_{n}^{L}}\left(\hat{x}_{n,p}\right)\left(\hat{M}\right)\big|\leq\frac{1}{L}.$
By the compactness of $\Omega^{L}$, passing the limit in $n$, for any $p\in\mathbb{N}$,
$\big|\hat{\mu}_{p}^{L}\left(\Omega^{L}\right)-\hat{\mu}^{L}_{p}\left(\hat{M}\right)\big|=\big|\hat{\mu}_{p}^{L}\left(\Omega^{L}\right)-\alpha^{L}\big|\leq\frac{1}{L}.$
Given $\varepsilon>0$, choose $L_{0}\in\mathbb{N}$ large enough such that for any $p\in\mathbb{N}$,
$\hat{\mu}_{p}^{L_{0}}\left(\Omega^{L_{0}}\right)\geq\alpha-\varepsilon.$
Recalling that $\hat{\mu}^{L}_{p}\xrightarrow{\text{$\quad L\quad$}}\hat{\mu}_{p}$ is an increasing convergence, for any $L\geq L_{0}$ and any $p\in\mathbb{N}$, we have
$\hat{\mu}_{p}^{L}\left(\Omega^{L_{0}}\right)\geq\hat{\mu}_{p}^{L_{0}}\left(\Omega^{L_{0}}\right)\geq\alpha-\varepsilon.$
By the compactness of $\Omega^{L_{0}}$, passing the limit in $L$, then in $p$, we have
$\hat{\mu}\left(\Omega^{L_{0}}\right)\geq\alpha-\varepsilon.$
By the arbitrariness of $\varepsilon$, we have
$\hat{\mu}\left(\bigcup_{L}\Omega^{L}\right)=\hat{\mu}\left(\hat{M}\right)=\alpha.$
Therefore the $\hat{f}-$invariant set $\bigcap_{i\in\mathbb{Z}}\hat{f}^{i}\left(\bigcup_{L}\Omega^{L}\right)$ has $\hat{\mu}$ full measure in which for any $\hat{x}$, by definition, there is an increasing sequence $0=m_{0}<m_{1}<m_{2}<\cdots$ such that
$\sum_{m=m_{i}}^{m_{i+1}-1}\rho\left(\hat{f}^{m}\left(\hat{x}\right)\right)\geq m_{i+1}-m_{i}.$
Hence for $\hat{\mu}-$almost every $\hat{x}$,
$\lim_{n\to+\infty}\frac{1}{n}\sum_{i=0}^{n-1}\rho\left(\hat{f}\left(\hat{x}\right)\right)=\limsup_{n\to+\infty}\frac{1}{n}\sum_{i=0}^{n-1}\rho\left(\hat{f}\left(\hat{x}\right)\right)\geq 1>0.$
This implies for the projection $\mu$-almost every $x$, $\lambda^{+}(f,x)>0$ and $\hat{\mu}$ is an unstable lift of $\mu$ (see Section 3.2). ∎
The following proposition provides an upper bound for the volume growth of a sequence of curves, which can be viewed as a prototype of Theorem 2.1
Recall the small constant $\varepsilon_{f}$ in Proposition 4.1.
Proposition 4.10.
Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact surface $M$ and let $\sigma_{n}$ be a sequence of $\varepsilon_{f}$-bounded $C^{r}$ curves. We have
$\limsup_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(\sigma_{n}\right)\right)\leq\max\left\{h_{\rm top}(f),\,\,\frac{\log\left\|Df\right\|}{r}\right\}+4\log\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)+C_{r}$
where $C_{r}$ is a universal constant depending only on $r$.
Remark.
We prove Proposition 4.10 by reformulating it in a convex combination form. More precisely, it suffices to show that there exists $\alpha\in[0,1]$ such that
$\limsup_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(\sigma_{n}\right)\right)\leq\alpha\cdot h_{\rm top}(f)+\left(1-\alpha\right)\cdot\frac{\log\left\|Df\right\|}{r}+4\log\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)+C_{r}.$
The parameter $\alpha$ arises from the relative proportions of different dynamical time regimes that interpolate between topological entropy and the Yomdin term. One can see from the proof that $\alpha$ corresponds, for some point $x$, to the proportion of geometric times (see Definition 11) after excluding the times spent near periodic sources (which are referred to as trapping times in the proof of Proposition 4.10).
Proof of Proposition 4.10.
Recall that the periodic sources are denoted by $z_{1},z_{2},z_{3},\cdots$ with the corresponding radii $r_{z_{i}}$ (See Definition 4.7).
Given $n\in\mathbb{N}$ and a point $x\in\sigma_{n}$, we define the trapping time of $x$ by
$t_{n}(x):=\max\left\{0<k\leq n:~\exists 1\leq s\leq n,\text{ s.t. }x\in B\left(z_{s},k,r_{z_{s}}\right)\right\}$
and set $t_{n}(x):=0$ if for any $1\leq s\leq n$,
$x\notin B\left(z_{s},r_{z_{s}}\right).$
Given $0\leq m\leq n$, define
$\sigma^{m}_{n}:=\{x\in\sigma_{n}:~t_{n}(x)=m\}$
and choose any $m_{n}$ (may not be unique) such that
${\rm Vol}\left(f^{n}\left(\sigma^{m_{n}}_{n}\right)\right)=\max_{0\leq m\leq n}{\rm Vol}\left(f^{n}\left(\sigma^{m}_{n}\right)\right).$
We define two partitions:
  • Given $n\in\mathbb{N}$ and $E_{n}\subset[0,n)$, define
    $\Lambda_{E_{n}}:=\left\{x\in\sigma^{m_{n}}_{n}:~E(\hat{x})\cap[0,n)=E_{n}\right\}$
    where, as previously defined, E(x^)E($\hat${x}) is the set of geometric times of $\hat{x}$. Let $\mathcal{E}_{n}$ be the partition of $\sigma^{m_{n}}_{n}$ with such elements $\Lambda_{E_{n}}$ for all possible $E_{n}$. Note that for the cardinality, we have
    $\mathcal{E}_{n}\leq 2^{n}.$
    We fix an element $\Lambda_{E_{n}}\in\mathcal{E}_{n}$ that achieves the maximal volume growth among all possible $\widetilde{E}_{n}$, i.e.,
    ${\rm Vol}\left(f^{n}\left(\Lambda_{E_{n}}\right)\right)=\max_{\widetilde{E}_{n}}{\rm Vol}\left(f^{n}\left(\Lambda_{\widetilde{E}_{n}}\right)\right).$
  • Given $p\in\mathbb{N}$ and a family of integers
    $\gamma:=\left\{\left(\gamma^{\prime}_{i},\,\hat{\gamma}^{\prime}_{i},\,\gamma^{\prime\prime}_{i},\,\hat{\gamma}^{\prime\prime}_{i}\right)\right\}_{i\in[0,n)},$
    define
    $\Lambda^{p,m_{n}}_{n,E_{n},\gamma}=\Big\{x\in\Lambda_{E_{n}}:~\,i\in[0,n),$ $\frac{1}{p}\left\lceil\log\left\|Df^{p}_{f^{i}(x)}\right\|\right\rceil=\gamma^{\prime}_{i},\,\frac{1}{p}\left\lceil\log\left\|Df^{p}\left(\hat{f}^{i}\left(\hat{x}\right)\right)\right\|\right\rceil=\hat{\gamma}^{\prime}_{i},\,$
    $\left\lceil\log\left\|Df_{f^{i}(x)}\right\|\right\rceil=\gamma^{\prime\prime}_{i},\,\left\lceil\log\left\|Df\left(\hat{f}^{i}\left(\hat{x}\right)\right)\right\|\right\rceil=\hat{\gamma}^{\prime\prime}_{i}\,\Big\}.$
    Let $\mathcal{E}_{n,E_{n}}^{p,m_{n}}$ be the partition of $\Lambda_{E_{n}}$ with such elements $\Lambda^{p,m_{n}}_{n,E_{n},\gamma}$ for all possible $\gamma$. We note two facts:
    • For the cardinality, we have
      $\mathcal{E}_{n,E_{n}}^{p,m_{n}}\leq\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)^{4n}.$
    • For any $L\in\mathbb{N}$, by definitions (see the definition of $O^{p,m_{n},L}_{n,E^{L}_{n},\beta}$ in front of Proposition 4.8),
      $\Lambda^{p,m_{n}}_{n,E_{n},\gamma}\subset O^{p,m_{n},L}_{n,E^{L}_{n},\beta}$
      where the corresponding $\beta=\beta(\gamma):=\left\{\left(\beta^{\prime}_{i},\,\beta^{\prime\prime}_{i}\right)\right\}_{i\in[0,n)}$ is defined by
      $\beta^{\prime}_{i}:=\gamma^{\prime}_{i},\,\,\beta^{\prime\prime}_{i}:=\hat{\gamma}^{\prime}_{i},\,\,i\in[m_{n},n)\cap E_{n}^{L}\,\,\text{ and }\,\,\beta^{\prime}_{i}:=\gamma^{\prime\prime}_{i},\,\,\beta^{\prime\prime}_{i}:=\hat{\gamma}^{\prime\prime}_{i},\,\,i\in[0,n)\setminus\left([m_{n},n)\cap E_{n}^{L}\right).$ (19)
For each $p$, we fix an element $\Lambda^{p,m_{n}}_{n,E_{n},\gamma}\in\mathcal{E}_{n,E_{n}}^{p,m_{n}}$ that achieves the maximal volume growth, i.e.,
${\rm Vol}\left(f^{n}\left(\Lambda^{p,m_{n}}_{n,E_{n},\gamma}\right)\right)=\max_{\widetilde{\gamma}}\,{\rm Vol}\left(f^{n}\left(\Lambda^{p,m_{n}}_{n,E_{n},\widetilde{\gamma}}\right)\right).$
We fix any point $x_{n,p}\in\Lambda^{p,m_{n}}_{n,E_{n},\gamma}$. Up to considering sub-sequences in $n,L,p$, we assume
$\delta_{[m_{n},n)\cap E_{n}^{L}}\left(\hat{x}_{n,p}\right)\xrightarrow{\text{$\quad n\quad$}}\hat{\mu}^{L}_{p},\quad\quad\hat{\mu}^{L}_{p}\xrightarrow{\text{$\quad L\quad$}}\hat{\mu}_{p},\quad\quad\hat{\mu}_{p}\xrightarrow{\text{$\quad p\quad$}}\hat{\mu}$ (20)
for some Borel measures (not necessarily probability measures) $\hat{\mu}^{L}_{p},\hat{\mu}_{p}$ and $\hat{\mu}$ (denote by $\mu^{L}_{p},\mu_{p}$ and $\mu$ the corresponding projections on $M$). By the convergences of these measures, we have
[…]
where $\alpha^{L},\alpha\in[0,1]$ are defined (independent of $p$) by
$\alpha^{L}:=\hat{\mu}^{L}_{p}(\hat{M})=\mu^{L}_{p}(M),\quad\alpha:=\hat{\mu}_{p}(\hat{M})=\hat{\mu}(\hat{M})=\mu_{p}(M)=\mu(M).$
By Proposition 4.8, for any $p,L\in\mathbb{N}$, recalling that $\gamma$ is determined by $\Lambda_{E_{n}}$ and $p$, and that $\beta$ is determined by $\gamma$ as (19),
$\log{\rm Vol}\left(f^{n}\left(\sigma_{n}\right)\right)$ $\leq\log\left(n+1\right)+\log{\rm Vol}\left(f^{n}\left(\sigma^{m_{n}}_{n}\right)\right)$ (21)
$\leq\log\left(n+1\right)+\log\#\mathcal{E}_{n}+\log{\rm Vol}\left(f^{n}\left(\Lambda_{E_{n}}\right)\right)$
$\leq\log\left(n+1\right)+\log\#\mathcal{E}_{n}+\log\#\mathcal{E}_{n,E_{n}}^{p,m_{n}}+\log{\rm Vol}\left(f^{n}\left(\Lambda^{p,m_{n}}_{n,E_{n},\gamma}\right)\right)$
$\leq\log\left(n+1\right)+\log\#\mathcal{E}_{n}+\log\#\mathcal{E}_{n,E_{n}}^{p,m_{n}}+\log{\rm Vol}\left(f^{n}\left(O^{p,m_{n},L}_{n,E^{L}_{n},\beta}\right)\right)$
$\leq\log\left(n+1\right)+n\log 2+4n\log\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)+\#\left([m_{n},n)\cap E^{L}_{n}\right)\cdot h_{\rm top}(f)$
$+\left(m_{n}+\#\left([m_{n},n)\setminus E^{L}_{n}\right)\right)\cdot\frac{\log\left\|Df\right\|}{r}+\sum_{k\in[m_{n},n)\cap E^{L}_{n}}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}+C_{n,L}$
where the error term $C_{n,L}$ satisfies
$\limsup_{L\to+\infty}\limsup_{n\to+\infty}\frac{C_{n,L}}{n}=C_{r}$
for some universal constant $C_{r}$ depending only on $r$.
Note that by definition, recalling $\rho(x,[v]):=\log\left\|Df_{x}\left(v\right)\right\|$,
$\frac{1}{n}\sum_{k\in[m_{n},n)\cap E^{L}_{n}}\frac{\beta^{\prime}_{k}-\beta^{\prime\prime}_{k}}{r-1}\leq\frac{1}{r-1}\int\frac{\log\left\|Df^{p}_{(\cdot)}\right\|-\sum_{i=0}^{p-1}\rho\circ\hat{f}^{i}}{p}d\,\left(\delta_{[m_{n},n)\cap E_{n}^{L}}\left(\hat{x}_{n,p}\right)\right)+10.$
Here the term ”1010” is introduced to compensate for the average discrepancy between the ceiling operation and the actual value.
Putting $\frac{1}{n}$ on both sides in (21), we have
$\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(\sigma_{n}\right)\right)\leq\alpha_{n,L}\cdot h_{\rm top}(f)+\left(1-\alpha_{n,L}\right)\cdot\frac{\log\left\|Df\right\|}{r}+\gamma_{n,L,p}$ (22)
where
$\alpha_{n,L}:=\frac{\#\left([m_{n},n)\cap E^{L}_{n}\right)}{n}$
and
[…]
$\quad+\frac{1}{r-1}\int\frac{\log\left\|Df^{p}_{(\cdot)}\right\|-\sum_{i=0}^{p-1}\rho\circ\hat{f}^{i}}{p}d\,\left(\delta_{[m_{n},n)\cap E_{n}^{L}}\left(\hat{x}_{n,p}\right)\right).$
We take $\limsup_{n\to+\infty}$, then take $\limsup_{L\to+\infty}$ on both sides in (22):
$\limsup_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(\sigma_{n}\right)\right)\leq\alpha\cdot h_{\rm top}(f)+\left(1-\alpha\right)\cdot\frac{\log\left\|Df\right\|}{r}+\limsup_{L\to+\infty}\limsup_{n\to+\infty}\gamma_{n,L,p}$
where
$\alpha:=\limsup_{L\to+\infty}\limsup_{n\to+\infty}\alpha_{n,L}.$
Consequently, for any $p\in\mathbb{N}$,
$\limsup_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(\sigma_{n}\right)\right)\leq\max\left\{h_{\rm top}(f),\,\,\frac{\log\left\|Df\right\|}{r}\right\}+\limsup_{L\to+\infty}\limsup_{n\to+\infty}\gamma_{n,L,p}.$
It remains to show the following claim.
Claim.
$\gamma_{n,L,p}$ satisfies
$\limsup_{p\to+\infty}\limsup_{L\to+\infty}\limsup_{n\to+\infty}\gamma_{n,L,p}\leq 4\log\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)+C_{r}$
for some universal constant $C_{r}$  [^34] The constant $C_{r}$ was defined earlier. We reuse the same notation here for simplicity. This causes no ambiguity, since the present $C_{r}$ may differ from the previous one only by a constant depending only on $r$. depending only on $r$.
Proof of Claim.
Let us first show the following two properties:
  • $\hat{\mu}_{p}$ and $\hat{\mu}$ defined in (20) are invariant measures,
  • $\hat{\mu}$ is the unique unstable lift of $\mu$, provided that $\mu$ is not zero measure.
To see this, by Proposition 4.9,
  • $\hat{\mu}_{p},\hat{\mu}$ are invariant measures,
  • $\mu$-almost every $x$ has at least one positive Lyapunov exponent (i.e., $\lambda^{+}(f,x)>0$),
  • $\hat{\mu}$ is an unstable lift of $\mu$.
We claim that $\mu$-almost every $x$ has a unique positive Lyapunov exponent (hence $\hat{\mu}$ is the unique unstable lift of $\mu$), due to the choice of $m_{n}$. Indeed, going by contradiction, if not, $\mu$ would have an ergodic component with two positive Lyapunov exponents and this ergodic component must support on the orbits of some periodic source $z$ with $\mu$ positive measure (see [29, Proposition 4.4] or [9, Fact in Claim 7.3]).
Let $r>0$ be small enough such that if $y\notin\cup_{i}B\left(f^{i}(z),r_{z}\right)$, then $f^{k}(y)\notin B(z,r)$ for all $k\in\mathbb{N}$. Indeed, it is enough to choose $r>0$ small enough such that for $0\leq j\leq T_{z}-1$ ($T_{z}$ denotes the period of $z$),
$f^{j}\left(B\left(z,r\right)\right)\subset B\left(f^{j}(z),r_{z}\right).$ (23)
To see this, we write $k=m\cdot T_{z}-j$ for some $0\leq j\leq T_{z}-1$. Since $y\notin B\left(f^{j}(z),r_{z}\right)$, by the choice of $r_{z}$ (see the second property in Definition 4.7),
$f^{j}\circ f^{k}(y)=f^{mT_{z}}(y)\notin B\left(f^{j}(z),r_{z}\right).$
By (23),
$f^{k}(y)\notin B\left(z,r\right).$
Since $\mu$ is not zero measure, we may assume for all large $n$,
$m_{n}+100<n.$
Since $x_{n,p}\in\sigma_{n}^{m_{n}}$, by the definition of $m_{n}$, there must exist some $0\leq i\leq m_{n}$ such that
$f^{i-1}(x_{n,p})\in B\left(f^{i-1}(z),r_{z}\right),\quad f^{i}(x_{n,p})\notin B\left(f^{i}(z),r_{z}\right).$
By the choice of $r_{z}$ (see the third property in Definition 4.7), we then have
$f^{i}(x_{n,p})\notin\bigcup_{i}B\left(f^{i}(z),r_{z}\right).$
As a consequence, for any $k\in[m_{n},n)\cap E_{n}^{L}$,
$f^{k}(x_{n,p})\notin B(z,r).$
This implies for large $n$,
$\delta_{[m_{n},n)\cap E_{n}^{L}}\left(x_{n,p}\right)\left(B(z,r)\right)=0$
which is a contradiction. Hence $\mu$-almost every $x$ must have a unique positive Lyapunov exponent and $\hat{\mu}$ is the unique unstable lift of $\mu$.
By taking limit in $n$, then in $L$, we have
$\quad\limsup_{L\to+\infty}\limsup_{n\to+\infty}\int\frac{\log\left\|Df^{p}_{(\cdot)}\right\|-\sum_{i=0}^{p-1}\rho\circ\hat{f}^{i}}{p}d\,\left(\delta_{[m_{n},n)\cap E_{n}^{L}}\left(\hat{x}_{n,p}\right)\right)$
$=\int\frac{\log\left\|Df^{p}_{(\cdot)}\right\|-\sum_{i=0}^{p-1}\rho\circ\hat{f}^{i}}{p}d\,\hat{\mu}_{p}.$
$=\int\frac{\log\left\|Df^{p}_{(\cdot)}\right\|}{p}d\,\mu_{p}-\int\rho d\,\hat{\mu}_{p}.$
Taking limit in $p$ on both sides, by Lemma 6.3, we then have
$\quad\limsup_{p\to+\infty}\limsup_{L\to+\infty}\limsup_{n\to+\infty}\int\frac{\log\left\|Df^{p}_{(\cdot)}\right\|-\sum_{i=0}^{p-1}\rho\circ\hat{f}^{i}}{p}d\,\left(\delta_{[m_{n},n)\cap E_{n}^{L}}\left(\hat{x}_{n,p}\right)\right)$
$\leq\int\limsup_{p\to+\infty}\frac{\log\left\|Df^{p}_{(\cdot)}\right\|}{p}d\,\mu-\int\rho d\,\hat{\mu}$.
Since $\hat{\mu}$ is the unique unstable lift of $\mu$, we have (see Section 3.2)
$\int\limsup_{p\to+\infty}\frac{\log\left\|Df^{p}_{(\cdot)}\right\|}{p}d\,\mu-\int\rho d\,\hat{\mu}=0.$
Hence
$\limsup_{p\to+\infty}\limsup_{L\to+\infty}\limsup_{n\to+\infty}\gamma_{n,L,p}\leq 4\log\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)+\log 2+10+C_{r}.$
We can put $\log 2+10$ into $C_{r}$ which completes the proof of the claim.
The proof of Proposition 4.10 is then complete. ∎
Now we use Proposition 4.10 to prove Theorem 2.1.
Proof of Theorem 2.1.
Write
$\chi(f):=\max\left\{h_{\rm top}(f),\,\,\frac{\log\left\|Df\right\|}{r}\right\}+4\log\left(\log\left\|Df\right\|+\log\left\|Df^{-1}\right\|+2\right)+C_{r},\qquad\alpha\in[0,1]$
where $C_{r}$ is the universal constant depending only $r$ from Proposition 4.10.
In the following arguments, the notation $\sup_{\sigma}$ refers to the supremum taken over all $C^{r}$ embedded curves $\sigma:[0,1]\to M$ with $\left\|\sigma\right\|_{C^{r}}\leq 1$ and $\left\|d_{t}\sigma\right\|\geq 1/2,\,t\in[0,1]$.
We claim
$\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}\left(f^{n}(\sigma)\right)\leq\chi(f).$ (24)
Assume by contradiction that
$\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}\left(f^{n}(\sigma)\right)>\chi(f).$
Then there exist $n_{k}\to+\infty$ and $C^{r}$ embedded curves $\sigma_{n_{k}}$ with $\left\|\sigma\right\|_{C^{r}}\leq 1$ and $\left\|d_{t}\sigma\right\|\geq 1/2,\,t\in[0,1]$ such that
$\limsup_{k\to+\infty}\frac{1}{n_{k}}\log{\rm Vol}\left(f^{n_{k}}(\sigma_{n_{k}})\right)>\chi(f).$
Since each $\sigma_{n_{k}}$ can be decomposed into a uniformly bounded number of $\varepsilon_{f}$-bounded curves (see Lemma 6.7), we may, without loss of generality, assume that the sequence $\{\sigma_{n_{k}}\}$ itself consists of $\varepsilon_{f}$-bounded curves. Hence, by Proposition 4.10,
$\limsup_{k\to+\infty}\frac{1}{n_{k}}\log{\rm Vol}\left(f^{n_{k}}(\sigma_{n_{k}})\right)\leq\chi(f),$
which is a contradiction.
Claim.
For any $q\geq 1$,
$\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}\left(f^{n}(\sigma)\right)=\limsup_{n\to+\infty}\frac{1}{qn}\log\sup_{\sigma}{\rm Vol}\left(f^{qn}(\sigma)\right).$
Proof of Claim.
Set
$a_{n}:=\sup_{\sigma}{\rm Vol}\left(f^{n}(\sigma)\right).$
The inequality $\geq$” follows since $\{qn\}$ is a sub-sequence of $\mathbb{N}$. For the reverse inequality, write $n=qk+m$ with $0\leq m<q$. Since $f^{m}$ has uniformly bounded $C^{1}$ norm for $0\leq m<q$, there exists $C>0$ such that
$a_{n}\leq C\,a_{qk}.$
Hence
$\frac{1}{n}\log a_{n}\leq\frac{qk}{n}\cdot\frac{1}{qk}\log a_{qk}+\frac{1}{n}\log C,$
and letting $n\to+\infty$ yields the result. ∎
Replacing $f$ by $f^{q}$ and putting $\frac{1}{q}$ on both sides in (24), by the arbitrariness of $q$, we get
$\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}\left(f^{n}(\sigma)\right)\leq\max\left\{h_{\rm top}(f),\,\,\frac{\lambda^{+}(f)}{r}\right\}.$
The proof of Theorem 2.1 is then complete. ∎
5 Volume growth of the tangent cocycle
In this section, we study the relationship between the volume growth of the tangent cocycle and entropy. More precisely, we establish the chain of inequalities:
$h_{\rm top}(f)\leq\liminf_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\leq\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\leq\max\left\{h_{\rm top}(f),\,\,\frac{\lambda^{+}(f)}{r}\right\}.$
The rightmost inequality is Theorem 2.2, while the leftmost inequality will be proved in Theorem 5.2.
5.1 Upper bound of the volume growth rate
The following lemma asserts the existence of a point $z$ at which the exponential growth rate of $\int_{M}\left\|Df^{n}_{x}\right\|$ can be detected on arbitrarily small neighborhoods of $z$.
Lemma 5.1.
Let $f$ be a $C^{1}$ diffeomorphism on a compact manifold $M$. There is a point $z$ such that for any neighborhood $U$ of $z$,
$\limsup_{n\to+\infty}\frac{1}{n}\log\int_{U}\left\|Df^{n}_{x}\right\|\,dx=\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx.$
Proof.
Write
$\lambda(U):=\limsup_{n\to+\infty}\frac{1}{n}\log\int_{U}\left\|Df^{n}_{x}\right\|\,dx.$
Suppose for contradiction that for every $z\in M$, there exists an open neighborhood $U_{z}$ such that $\lambda(U_{z})<\lambda(M)$.
By the compactness of $M$, there exists a finite sub-cover $\{U_{z_{i}}\}_{i=1}^{k}$ such that $M=\bigcup_{i=1}^{k}U_{z_{i}}$. Using the property that for non-negative sequences $a_{n,i}$,
$\limsup_{n\to+\infty}\frac{1}{n}\log\left(\sum_{i=1}^{k}a_{n,i}\right)=\max_{1\leq i\leq k}\left(\limsup_{n\to+\infty}\frac{1}{n}\log a_{n,i}\right),$
we have:
$\lambda(M)=\lambda\left(\bigcup_{i=1}^{k}U_{z_{i}}\right)=\max_{1\leq i\leq k}\lambda(U_{z_{i}})<\lambda(M)$
which is a contradiction. ∎
The main idea to prove Theorem 2.2 is to reduce the estimation of the volume growth of the tangent cocycle to that of one-dimensional submanifolds via Fubini’s theorem, which allows the two-dimensional integral of $\left\|Df^{n}_{x}\right\|$ to be decomposed into integrals along horizontal and vertical lines. The conclusion then follows by applying the curve volume growth estimate in Theorem 2.1.
Proof of Theorem 2.2.
Let $\mathcal{A}:=\left\{\left(U_{i},\varphi_{i}\right):~\varphi_{i}:U_{i}\to[0,1]^{2}\right\}_{0\leq i<N}$ be a finite family of charts that cover $M$. Let $z$ be the point in Lemma 5.1. Without loss of generality, we assume $z\in U_{0}$ and $U_{0}$ is a ball at $z$ with sufficiently small diameter such that for any line $\gamma$ in $0,1]^{2}$, $\varphi^{-1}_{0}\left(\gamma\right)$ is contained in some $C^{r}$ embedded curve $\sigma:[0,1]\to M$ with $\left\|\sigma\right\|_{C^{r}}\leq 1$ and $\left\|d_{t}\sigma\right\|\geq 1/2,\,t\in[0,1]$. Let $K$ be a sufficiently large number (independent of $n$) such that
  • for any $0\leq i<N$ and any curve $\sigma$ on $U_{i}$,
    $K^{-1}\cdot{\rm Vol}\left(\varphi_{i}\circ\sigma\right)\leq{\rm Vol}\left(\sigma\right)\leq K\cdot{\rm Vol}\left(\varphi_{i}\circ\sigma\right).$
  • for any $0\leq i<N$ and any $n\geq 1$,
    $\int_{U_{0}^{i}}\left\|Df^{n}_{x}\right\|\,dx\leq K\int_{\varphi_{0}\left(U_{0}^{i}\right)}\left\|DF^{n}_{\left(x_{1},x_{2}\right)}\right\|\,dx_{1}dx_{2}$ (25)
    where $\left(x_{1},x_{2}\right)$ denotes the standard coordinates in $\mathbb{R}^{2}$ and
    $F^{n}:=\varphi_{i}\circ f^{n}\circ\varphi^{-1}_{0},\quad U_{0}^{i}:=\left\{x\in U_{0}:~f^{n}\left(x\right)\in U_{i}\right\}.$
Given $n\in\mathbb{N}$, we fix any $0\leq i<N$ such that
$\int_{U_{0}}\left\|Df^{n}_{x}\right\|\,d\leq N\cdot\int_{U_{0}^{i}}\left\|Df^{n}_{x}\right\|\,dx.$ (26)
Given a $C^{1}$ map $G:=(G_{1},G_{2})$ on $\mathbb{R}^{2}$, denote partial derivative vectors of $G$ by
$\partial_{1}G_{(x_{1},x_{2})}:=\left(\frac{\partial G_{1}}{\partial x_{1}}\left(x_{1},x_{2}\right),\frac{\partial G_{2}}{\partial x_{1}}\left(x_{1},x_{2}\right)\right),\quad\partial_{2}G_{(x_{1},x_{2})}:=\left(\frac{\partial G_{1}}{\partial x_{2}}\left(x_{1},x_{2}\right),\frac{\partial G_{2}}{\partial x_{2}}\left(x_{1},x_{2}\right)\right).$
Note that, by standard linear algebra,
$\left\|DG_{\left(x_{1},x_{2}\right)}\right\|\leq\left\|\partial_{1}G_{(x_{1},x_{2})}\right\|+\left\|\partial_{2}G_{(x_{1},x_{2})}\right\|$
where $\left\|\cdot\right\|$ denotes the usual operator norm for matrices and the usual Euclidean norm for vectors. Consider the foliation of horizontal lines
$\left\{[0,1]\times x_{2}:\ x_{2}\in[0,1]\right\}.$
By Fubini’s theorem and by the choice of $K$, we have
$\int_{\varphi_{0}\left(U_{0}^{i}\right)}\left\|\partial_{1}F^{n}_{\left(x_{1},x_{2}\right)}\right\|\,dx_{1}dx_{2}$ $=\int_{0}^{1}\left(\int_{\varphi_{0}\left(U_{0}^{i}\right)\cap([0,1]\times x_{2})}\left\|\partial_{1}F^{n}_{\left(x_{1},x_{2}\right)}\right\|\,dx_{1}\right)dx_{2}$
$=\int_{0}^{1}{\rm Vol}\left(F^{n}\left(\varphi_{0}\left(U_{0}^{i}\right)\cap\left([0,1]\times x_{2}\right)\right)\right)dx_{2}$
$\leq\max_{0\leq x_{2}\leq 1}{\rm Vol}\left(F^{n}\left(\varphi_{0}\left(U_{0}^{i}\right)\cap\left([0,1]\times x_{2}\right)\right)\right)$
$\leq K\cdot\max_{0\leq x_{2}\leq 1}{\rm Vol}\left(f^{n}\left(U_{0}^{i}\cap\varphi_{0}^{-1}\left([0,1]\times x_{2}\right)\right)\right)$
$\leq K\cdot\max_{0\leq x_{2}\leq 1}{\rm Vol}\left(f^{n}\left(\varphi_{0}^{-1}\left([0,1]\times x_{2}\right)\right)\right).$
Similarly, by considering vertical lines,
$\int_{\varphi_{0}\left(U_{0}^{i}\right)}\left\|\partial_{2}F^{n}_{\left(x_{1},x_{2}\right)}\right\|\,dx_{1}dx_{2}\leq K\cdot\max_{0\leq x_{1}\leq 1}{\rm Vol}\left(f^{n}\left(\varphi_{0}^{-1}\left(x_{1}\times[0,1]\right)\right)\right).$
We define $\sigma_{n}$ to be the pre-image under $\varphi_{0}$ of a horizontal or vertical line in $0,1]^{2}$ whose image under $f^{n}$ attains the maximal volume among all such coordinate lines above, namely,
${\rm Vol}\left(f^{n}\left(\sigma_{n}\right)\right)=\max_{0\leq x_{1},x_{2}\leq 1}\left\{{\rm Vol}\left(f^{n}\left(\varphi_{0}^{-1}\left([0,1]\times x_{2}\right)\right)\right),\,{\rm Vol}\left(f^{n}\left(\varphi_{0}^{-1}\left(x_{1}\times[0,1]\right)\right)\right)\right\}.$
Hence
$\int_{\varphi_{0}\left(U_{0}^{i}\right)}\left\|DF^{n}_{(x_{1},x_{2})}\right\|\,dx_{1}dx_{2}$ $\leq\int_{\varphi_{0}\left(U_{0}^{i}\right)}\left(\left\|\partial_{1}F^{n}_{(x_{1},x_{2})}\right\|+\left\|\partial_{2}F^{n}_{(x_{1},x_{2})}\right\|\right)\,dx_{1}dx_{2}$
$\leq 2K\cdot{\rm Vol}\left(f^{n}\left(\sigma_{n}\right)\right).$
Applying Theorem 2.1 and recalling (25) and (26) , we have
$\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx$ $=\limsup_{n\to+\infty}\frac{1}{n}\log\int_{U_{0}}\left\|Df^{n}_{x}\right\|\,dx$
$\leq\limsup_{n\to+\infty}\frac{1}{n}\log{\rm Vol}\left(f^{n}\left(\sigma_{n}\right)\right)$
$\leq\max\left\{h_{\rm top}(f),\,\,\frac{\lambda^{+}(f)}{r}\right\}.$
The proof of Theorem 2.2 is then complete. ∎
5.2 Lower bound of the volume growth rate
The goal of this section is to establish the following result, which can be viewed as a consequence of Pesin theory and is, in a certain sense, already known. The distinction here lies in the formulation: Theorem 5.2 is stated directly in terms of the derivative $Df^{n}_{x}$, rather than the induced map $(Df^{n}_{x})^{\wedge}$ on exterior algebras of tangent spaces, as considered by Przytycki [30] and Kozlovski [23].
Theorem 5.2.
Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact surface $M$.
$h_{\rm top}(f)\leq\liminf_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx.$
We first recall a result of Katok (Theorem 1.1 in [20]).
Let $\mu$ be an ergodic measure. Given $\lambda\in(0,1)$, we define $S_{\lambda}(n,\varepsilon)\subset M$ to be a finite set such that
$\mu\left(\bigcup_{x\in S_{\lambda}(n,\varepsilon)}B(x,n,\varepsilon)\right)\geq\lambda,$
and whose cardinality is minimal among all finite subsets of $M$ with this property. Given a subset $\Omega$, recall (see [33] for background in ergodic theory) that a subset $S\subset\Omega$ is called a $(n,\varepsilon)$-separated set if for any $x,y\in S$ with $x\neq y$, there is $0\leq k<n$ such that $d(f^{k}(x),f^{k}(y))>\varepsilon$. Moreover, a $(n,\varepsilon)$-separated set $S$ is called maximal if its cardinality is larger than or equal to the cardinality of any other $(n,\varepsilon)$-separated set. Given a subset $\Omega$, let $S(n,\varepsilon,\Omega)$ denotes a maximal $(n,\varepsilon)$-separated set of $\Omega$. By maximality, the union of the dynamical balls $\{B(x,n,\varepsilon)\}_{x\in S(n,\varepsilon,\Omega)}$ covers $\Omega$.
Lemma 5.3 (Katok, [20]).
Let $f$ be a homeomorphism on a compact metric space $X$ and let $\mu$ be an ergodic measure. Then for any $\lambda\in(0,1)$,
$h(f,\mu)=\lim_{\varepsilon\to 0}\liminf_{n\to+\infty}\frac{1}{n}\log\#S_{\lambda}(n,\varepsilon)=\lim_{\varepsilon\to 0}\limsup_{n\to+\infty}\frac{1}{n}\log\#S_{\lambda}(n,\varepsilon).$
The following lemma, due to Przytycki, as an application of Pesin theory, provides a uniform lower bound on the integral of the derivative over dynamical balls centered at typical points w.r.t. a hyperbolic ergodic measure.
Lemma 5.4.
[30, Lemma 2.1] Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact surface $M$ and let $\mu$ be a hyperbolic ergodic measure. There are a subset $\Omega\subset M$ with $\mu\left(\Omega\right)>\frac{1}{2}$ and a constant $C>0$ such that for any $z\in\Omega$, any $n\in\mathbb{N}$ and any $0<\varepsilon<2026$,
$\int_{B(z,n,\varepsilon)}\left\|Df^{n}_{x}\right\|\,dx\geq C\cdot\varepsilon^{2}\cdot\frac{1}{2^{n}}.$
Remark 5.5.
The original statement of Lemma 5.4 is formulated for the induced map $\left(Df^{n}_{x}\right)^{\wedge}$ acting on the exterior algebra of the tangent space. However, one can see from the proof that the same conclusion holds when $\left(Df^{n}_{x}\right)^{\wedge}$ is replaced by $\left\|Df^{n}_{x}\right\|$, since for a hyperbolic measure on a surface the unstable dimension is equal to one.
Proof of Theorem 5.2.
Given a hyperbolic ergodic measure $\mu$, by Lemma 5.4, there are a subset $\Omega\subset M$ with $\mu\left(\Omega\right)\geq\frac{1}{2}$ and a constant $C>0$ such that for any $z\in\Omega$, any $n\in\mathbb{N}$ and any $0<\varepsilon<2026$,
$\int_{B(z,n,\varepsilon)}\left\|Df^{n}_{x}\right\|\,dx\geq C\cdot\varepsilon^{2}\cdot\frac{1}{2^{n}}.$
Let $S(n,2\varepsilon,\Omega)$ be a maximal $(n,2\varepsilon)$ separated set of $\Omega$. By definition, for any $z_{1},z_{2}\in S(n,2\varepsilon,\Omega)$,
$B(z_{1},n,\varepsilon)\cap B(z_{2},n,\varepsilon)=\emptyset.$
We have
$\int_{M}\left\|Df^{n}_{x}\right\|\,dx$ $\geq\#S(n,2\varepsilon,\Omega)\cdot\min_{z\in S(n,2\varepsilon,\Omega)}\int_{B(z,n,\varepsilon)}\left\|Df^{n}_{x}\right\|\,dx$
$\geq\#S(n,2\varepsilon,\Omega)\cdot C\cdot\varepsilon^{2}\cdot\frac{1}{2^{n}}.$
Then
$\liminf_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\geq\liminf_{n\to+\infty}\frac{1}{n}\log\#S(n,2\varepsilon,\Omega)-\log 2.$
Since $\mu(\Omega)>\frac{1}{2}$, we have $S(n,2\varepsilon,\Omega)\geq\#S_{\frac{1}{2}}(n,2\varepsilon)$. By Lemma 5.3 and the arbitrariness of $\varepsilon$, we then have
$\liminf_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\geq h(f,\mu)-\log 2.$
By the variational principle, we have
$\liminf_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\geq h_{\rm top}(f)-\log 2.$ (27)
Note that
$\liminf_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx=\liminf_{n\to+\infty}\frac{1}{qn}\log\int_{M}\left\|Df^{qn}_{x}\right\|\,dx,\quad h_{\rm top}(f)=\frac{h_{\rm top}(f^{q})}{q}.$
Replacing $f$ by $f^{q}$ and putting $\frac{1}{q}$ on both sides in (27), by the arbitrariness of $q$, we finally get the conclusion:
$\liminf_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\geq h_{\rm top}(f).$
6 Appendix
Lemma 6.1.
Let $f$ be a homeomorphism on a compact metric space $(X,d)$. Given $n\in\mathbb{N}$ and $\varepsilon>0$, let $\left\{B(x_{i},n,\varepsilon)\right\}_{1\leq i\leq N}$ be a collection of mutually disjoint $(n,\varepsilon)$ dynamical balls. There is $0\leq k_{n}\leq n-1$ such that
$\sum_{i=1}^{N-1}d\left(f^{k_{n}}\left(x_{i}\right),f^{k_{n}}\left(x_{i+1}\right)\right)\geq\frac{\left(N-1\right)\varepsilon}{n}.$
Proof.
Since the dynamical balls $\left\{B(x_{i},n,\varepsilon)\right\}_{1\leq i\leq N}$ are mutually disjoint,
$\max_{0\leq k\leq n-1}d\left(f^{k}(x_{i}),f^{k}(x_{i+1})\right)\geq\varepsilon.$
Hence
$\sum_{i=1}^{N-1}\sum_{k=0}^{n-1}d\left(f^{k}(x_{i}),f^{k}(x_{i+1})\right)\geq\sum_{i=1}^{N-1}\max_{0\leq k\leq n-1}d\left(f^{k}(x_{i}),f^{k}(x_{i+1})\right)\geq(N-1)\varepsilon.$
Changing the order of summation yields:
$\sum_{k=0}^{n-1}\sum_{i=1}^{N-1}d\left(f^{k}(x_{i}),f^{k}(x_{i+1})\right)\geq(N-1)\varepsilon.$
Hence, there must exist at least one index $k_{n}\in\{0,1,\dots,n-1\}$ such that
$\sum_{i=1}^{N-1}d\left(f^{k_{n}}\left(x_{i}\right),f^{k_{n}}\left(x_{i+1}\right)\right)\geq\frac{\left(N-1\right)\varepsilon}{n}.$
Recall that $h_{\rm top}(f,\Omega)$ denotes the topological entropy on a subset $\Omega$, defined in the usual way by
$h_{\rm top}\left(f,\Omega\right):=\lim_{\varepsilon\to 0}\limsup_{n\to+\infty}\frac{1}{n}\log\#S(n,\varepsilon,\Omega)$
where $S(n,\varepsilon,\Omega)$ denotes a maximal $(n,\varepsilon)$-separated subset of $\Omega$.
Lemma 6.2.
Let $f$ be a $C^{1}$ diffeomorphism on a compact manifold $M$ and let $\sigma:[0,1]\to M$ be a $C^{1}$ curve. We have
$h_{\rm top}\left(f,\sigma\right)\leq\limsup_{n\to+\infty}\frac{1}{n}\log^{+}{\rm Vol}\left(f^{n}\left(\sigma\right)\right).$
Proof.
We may assume $h_{\rm top}\left(f,\sigma\right)>0$, otherwise the inequality is trivially true.
Given $0<\delta<h_{\rm top}\left(f,\sigma\right)/2$, let $\varepsilon>0$ be such that
$\limsup_{n\to+\infty}\frac{1}{n}\log\#S(n,2\varepsilon,\sigma)\geq h_{\rm top}\left(f,\sigma\right)-\delta>0$
where $S(n,2\varepsilon,\sigma)$ denotes a maximal $(n,2\varepsilon)$-separated set of $\sigma$. Write $S_{n}:=\#S(n,2\varepsilon,\sigma)$ for simplicity. We enumerate the points in $S(n,2\varepsilon,\sigma)$ as $\{x_{1},x_{2},\cdots x_{S_{n}}\}$ in such a way that $\sigma^{-1}\left(x_{i}\right)<\sigma^{-1}\left(x_{i+1}\right)$. Note that By definition, for any $z,z^{\prime}\in S(n,2\varepsilon,\sigma)$,
$B(z,n,\varepsilon)\cap B(z^{\prime},n,\varepsilon)=\emptyset.$
By Lemma 6.1, there is $0\leq k_{n}\leq n-1$ such that
$\sum_{i=1}^{S_{n}-1}d\left(f^{k_{n}}\left(x_{i}\right),f^{k_{n}}\left(x_{i+1}\right)\right)\geq\frac{\left(S_{n}-1\right)\varepsilon}{n}.$
As a consequence,
${\rm Vol}\left(f^{k_{n}}\left(\sigma\right)\right)\geq\frac{\left(S_{n}-1\right)\varepsilon}{n}.$
Hence, for all large $n$,
$\frac{1}{k_{n}}\log^{+}{\rm Vol}\left(f^{k_{n}}\left(\sigma\right)\right)\geq\frac{n}{k_{n}}\cdot\left(\frac{1}{n}\log\frac{\left(S_{n}-1\right)\varepsilon}{n}\right)\geq\frac{1}{n}\log\frac{\left(S_{n}-1\right)\varepsilon}{n}.$
Note that $k_{n}\to+\infty$ as $n\to+\infty$. Putting $\limsup_{n\to+\infty}$” on both sides, we then have
$\limsup_{n\to+\infty}\frac{1}{n}\log^{+}{\rm Vol}\left(f^{n}\left(\sigma\right)\right)\geq\limsup_{n\to+\infty}\frac{1}{k_{n}}\log^{+}{\rm Vol}\left(f^{k_{n}}\left(\sigma\right)\right)\geq h_{\rm top}\left(f,\sigma\right)-\delta.$
By the arbitrariness of $\delta$, we then have
$\limsup_{n\to+\infty}\frac{1}{n}\log^{+}{\rm Vol}\left(f^{n}\left(\sigma\right)\right)\geq h_{\rm top}\left(f,\sigma\right).$
The following lemma establishes an upper semi-continuity property for integrals of sub-additive potentials along converging sequences of invariant measures. This result is a special case of [12, Lemma 2.3].
Lemma 6.3.
Let $f$ be a homeomorphism on a compact metric space $X$. Let $\mu_{n}$ be a sequence of invariant probability measures converging to some invariant probability measure $\mu$ and let $\varphi_{n}$ be a sequence of sub-additive continuous functions. We have
$\limsup_{n\to+\infty}\int\frac{\varphi_{n}}{n}d\,\mu_{n}\leq\int\limsup_{n\to+\infty}\frac{\varphi_{n}}{n}d\,\mu.$
Proof.
Fix $m\geq 1$. For any $n$, write $n=km+r$ with $0\leq r<m$. By the sub-additivity of $\varphi_{n}$, we have
$\varphi_{n}\leq\sum_{j=0}^{k-1}\varphi_{m}\circ f^{jm}+\varphi_{r}\circ f^{km}.$
Integrating w.r.t. the $f$-invariant measure $\mu_{n}$, we obtain
$\int\varphi_{n}\,d\mu_{n}\leq k\int\varphi_{m}\,d\mu_{n}+C_{m},$
where $C_{m}:=\max_{0\leq r<m}\max_{x}\varphi_{r}(x)$. Dividing by $n$ yields
$\int\frac{\varphi_{n}}{n}\,d\mu_{n}\leq\frac{k}{n}\int\varphi_{m}\,d\mu_{n}+\frac{C_{m}}{n}\leq\frac{1}{m}\int\varphi_{m}\,d\mu_{n}+\frac{C_{m}}{n}.$
Since $\mu_{n}\to\mu$ and $\varphi_{m}$ is continuous, letting $n\to+\infty$ gives
$\limsup_{n\to+\infty}\int\frac{\varphi_{n}}{n}\,d\mu_{n}\leq\frac{1}{m}\int\varphi_{m}\,d\mu.$
Taking the infimum over $m\geq 1$, and applying Kingman’s Sub-additive Ergodic Theorem (specifically, $\inf_{m}\frac{1}{m}\int\varphi_{m}\,d\mu=\int\lim_{n}\frac{\varphi_{n}}{n}\,d\mu$), we conclude
$\limsup_{n\to+\infty}\int\frac{\varphi_{n}}{n}\,d\mu_{n}\leq\inf_{m\geq 1}\frac{1}{m}\int\varphi_{m}\,d\mu=\int\limsup_{n\to+\infty}\frac{\varphi_{n}}{n}\,d\mu.$
Lemma 6.4 (Reverse Fatou Lemma).
Let $X$ be a compact metric space and let $\mu$ be a Borel probability measure on $X$. Let $\{\varphi_{n}\}_{n\geq 1}$ be a sequence of measurable functions on $X$ such that $\sup_{n\geq 1}\left\|\varphi_{n}\right\|_{\infty}<+\infty$. Then
$\int_{X}\limsup_{n\to+\infty}\varphi_{n}\,d\mu\geq\limsup_{n\to+\infty}\int_{X}\varphi_{n}\,d\mu.$
The following lemma (see the book [28, Lemma 3.33]) gives an abstract comparison between the pointwise exponential growth rate and the exponential growth rate of the integral. We provide a proof for completeness.
Lemma 6.5.
Let $(X,\mathcal{A},\mu)$ be a probability space and $\varphi_{n}:X\to(0,+\infty)$ a sequence of integrable functions. Then
$\limsup_{n\to+\infty}\frac{1}{n}\log\varphi_{n}(x)\leq\limsup_{n\to+\infty}\frac{1}{n}\log\int_{X}\varphi_{n}\,d\mu,$
for $\mu$-a.e. $x\in X$.
Proof.
Write
$\lambda:=\limsup_{n\to+\infty}\frac{1}{n}\log\int_{X}\varphi_{n}\,d\mu.$
Given $\varepsilon>0$, we have
$\int_{X}\varphi_{n}\,d\mu\geq\int_{A_{n}}\varphi_{n}\,d\mu\geq\mu(A_{n})\cdot e^{n(\lambda+\varepsilon)}$
where
$A_{n}:=\left\{x\in X:~\varphi_{n}(x)\geq e^{n(\lambda+\varepsilon)}\right\}.$
Rearranging and taking the limit yields:
$\limsup_{n\to+\infty}\frac{1}{n}\log\mu(A_{n})\leq\limsup_{n\to+\infty}\frac{1}{n}\log\int_{X}\varphi_{n}\,d\mu-(\lambda+\varepsilon)=-\varepsilon<0.$
Hence
$\sum_{n=1}^{+\infty}\mu(A_{n})<+\infty.$
By the Borel-Cantelli Lemma, for $\mu$-a.e. $x$, there exists $N(x)$ such that for all $n\geq N(x)$,
$\varphi_{n}(x)<e^{n(\lambda+\varepsilon)},$
which proves the lemma by the arbitrariness of $\varepsilon$. ∎
Lemma 6.6.
Let $f$ be a homeomorphism on a compact metric space $X$. Let $\mu$ be an ergodic probability measure and let $A$ be a subset with $\mu(A)>0$. Let $\varphi:X\to\mathbb{R}$ be a measurable function. If for any $x\in A$, there exists $n_{x}\in\mathbb{N}$ such that $\varphi(f^{-n}(x))=0$ for all $n\geq n_{x}$, then $\varphi(x)=0$ for $\mu$-a.e. $x$.
Proof.
Define
$Z:=\left\{x\in X\mid\exists N\in\mathbb{N}\text{ s.t. }\forall n\geq N,\varphi(f^{-n}(x))=0\right\}.$
By definition, $Z$ is $f$-invariant, and $A\subset Z$. Since $\mu(A)>0$, ergodicity implies $\mu(Z)=1$.
If $\mu(\{\varphi\neq 0\})>0$, the Poincaré recurrence theorem applied to $f^{-1}$ implies that for $\mu$-a.e. $x\in\{\varphi\neq 0\}$, the backward orbit $f^{-n}(x)$ returns to $\{\varphi\neq 0\}$ infinitely often. This contradicts $\mu(Z)=1$. Thus, $\varphi=0$ $\mu$-a.e. ∎
Lemma 6.7 (Decomposition into $\varepsilon$-bounded curves).
Let $M$ be a compact manifold and let $r>1$. Then for any $0<\varepsilon\leq\frac{1}{100}$, any $C^{r}$ curve $\sigma:[0,1]\to M$ with
$\left\|\sigma\right\|_{C^{r}}=\max_{1\leq s\leq r}\left\|d^{s}\sigma\right\|\leq 1\,\text{ and }\,\left\|d_{t}\sigma\right\|\geq 1/2,\,t\in[0,1],$
there are $\varepsilon$-bounded $C^{r}$ curves $\{\sigma_{j}\}_{1\leq j\leq\left\lceil\frac{1}{\varepsilon}\right\rceil}$ such that
$\sigma\left([0,1]\right)=\bigcup_{j=1}^{\left\lceil\frac{1}{\varepsilon}\right\rceil}\sigma_{j}\left([0,1]\right).$
Proof.
Take $N=\left\lceil\frac{1}{\varepsilon}\right\rceil$ and define the step size $\delta=\frac{1}{N}$. Note that by assumption $\varepsilon\leq\frac{1}{100}$, we have $\delta\leq\varepsilon\leq\frac{1}{100}$.
Define the translation constants $b_{j}$ for $j=1,\dots,N$ as follows:
$b_{j}=\begin{cases}(j-1)\delta,&1\leq j<N\\ 1-\delta,&j=N\end{cases}$
The affine maps $\theta_{j}:[0,1]\to[0,1]$ are given by $\theta_{j}(t)=\delta t+b_{j}$. By construction, $\theta_{j}([0,1])=[b_{j},b_{j}+\delta]\subset[0,1]$, and the union of these images covers $[0,1]$ (with the last interval $1-\delta,1]$ potentially overlapping with the previous one).
Let $\sigma_{j}=\sigma\circ\theta_{j}$. We verify the $\varepsilon$-boundedness for each $1\leq j\leq N$:
  1. 1.
    First derivative bound.
    Since $\left\|d_{t}\sigma\right\|\leq 1$ and $\theta_{j}^{\prime}=\delta$, we have
    $\left\|d\sigma_{j}\right\|=\sup_{t\in[0,1]}\left\|d_{t}\sigma\cdot\theta_{j}^{\prime}\right\|\leq 1\cdot\delta\leq\varepsilon.$
  2. 2.
    Higher order derivative bound.
    For $2\leq s\leq r$, the chain rule gives $\left\|d^{s}\sigma_{j}\right\|\leq\delta^{s}\left\|d^{s}\sigma\right\|\leq\delta^{s}$. Recalling the lower bound $\left\|d_{t}\sigma\right\|\geq 1/2$, we have $\left\|d\sigma_{j}\right\|\geq\frac{1}{2}\delta$. Consequently,
    $\frac{\left\|d^{s}\sigma_{j}\right\|}{\left\|d\sigma_{j}\right\|}\leq\frac{\delta^{s}}{\frac{1}{2}\delta}=2\delta^{s-1}\leq 2\delta\leq\frac{2}{100}<\frac{1}{6},$
    which satisfies the boundedness condition for all $2\leq s\leq r$.
Therefore, each $\sigma_{j}$ is $\varepsilon$-bounded, and $\sigma([0,1])=\bigcup_{j=1}^{N}\sigma_{j}([0,1])$. ∎
The following example shows that the local volume growth rate can still be positive in a small neighborhood of a repeller. Here the positivity of the local volume growth is a purely local phenomenon: although the curve is exponentially restricted to an increasingly small neighborhood of the repeller, the rapidly growing oscillation frequency produces a strong transverse amplification of derivatives. The resulting growth reflects a subtle balance between geometric restriction (uniform expansion) and oscillatory stretching.
Example 6.8.
Let $A=\operatorname{diag}(a,3)$ with $a>1$, and let $\sigma:[0,1]\to\mathbb{R}^{2}$ be a $C^{2}$ curve defined by $\sigma(x)=(x,x^{5}\sin(1/x))$. The volume (i.e., arc length) growth rate restricted in $1,1]^{2}$  [^38] The restriction to $1,1]^{2}$ is meant to mimic the restriction to a dynamical ball, which is justified here by the uniform expansion of the map., defined by
${\rm Vol}(A,\sigma)_{[-1,1]^{2}}:=\limsup_{n\to+\infty}\frac{1}{n}\log\operatorname{Length}(A^{n}\sigma\cap[-1,1]^{2}),$
is given by:
${\rm Vol}(A,\sigma)_{[-1,1]^{2}}=\begin{cases}\frac{1}{5}\log 3,&1<a\leq 3^{1/5}\\ \log 3-4\log a,&3^{1/5}<a<3^{1/4}\\ 0,&a\geq 3^{1/4}\end{cases}$
In particular, the growth rate is positive if and only if $1<a<3^{1/4}$.
Proof.
Step 1. Parametrization and derivatives.
Inside $1,1]^{2}$, the image curve $A^{n}\sigma$ admits the parametrization
$\gamma_{n}(u)=(u,g_{n}(u)),\qquad u\in[0,1],$
where
$g_{n}(u)=3^{n}a^{-5n}u^{5}\sin\!\left(\frac{a^{n}}{u}\right).$
The admissible domain to have $\gamma_{n}(u)=(u,g_{n}(u))\in[-1,1]^{2}$ is determined by
$u\in[0,1],\qquad|g_{n}(u)|\leq 1.$
Denote by $I_{n}$ the set of $u$ satisfying these conditions. The length of $A^{n}\sigma\cap[-1,1]^{2}$ is
$L_{n}:=\int_{I_{n}}\sqrt{1+|g_{n}^{\prime}(u)|^{2}}\,du.$
A direct computation gives
$g_{n}^{\prime}(u)=5\cdot 3^{n}a^{-5n}u^{4}\sin\!\left(\frac{a^{n}}{u}\right)-3^{n}a^{-4n}u^{3}\cos\!\left(\frac{a^{n}}{u}\right).$
Step 2. Case analysis.
The figure 3 below illustrates how the local volume growth transitions from bounded oscillations to amplified and eventually explosive behavior as the parameter $a$ varies across different regimes.
Figure 3: Local volume growth for different value of $a$
  • Case 1: $a\geq 3^{1/4}$.
    In this range, $3a^{-4}\leq 1$. For $u\in[0,1]$, we have
    $|g_{n}(u)|\leq(3a^{-5})^{n}\leq 1,$
    which implies $I_{n}=[0,1]$. Also note
    $|g_{n}^{\prime}(u)|\leq 5\cdot 3^{n}a^{-5n}+3^{n}a^{-4n}=5(3a^{-5})^{n}+(3a^{-4})^{n}.$
    As $a>1$ and $3a^{-4}\leq 1$, there exists $C>0$ such that $|g_{n}^{\prime}(u)|\leq C$ for all $n$ and $u\in I_{n}$. Hence
    $1\leq\sqrt{1+|g_{n}^{\prime}(u)|^{2}}\leq\sqrt{1+C^{2}},$
    which implies
    ${\rm Vol}(A,\sigma)_{[-1,1]^{2}}=\limsup_{n\to+\infty}\frac{1}{n}\log L_{n}\leq 0.$
    Since $I_{n}=[0,1]$, the image curve has length at least one. Hence
    ${\rm Vol}(A,\sigma)_{[-1,1]^{2}}=0$
  • Case 2: $3^{1/5}<a<3^{1/4}$.
    In this range, $3a^{-5}<1$ and $3a^{-4}>1$. Since
    $|g_{n}(u)|\leq(3a^{-5})^{n}\leq 1,$
    we have $I_{n}=[0,1]$ as in the previous case.
    Note
    $5\cdot 3^{n}a^{-5n}u^{4}\sin\!\left(\frac{a^{n}}{u}\right)\leq 5\cdot 3^{n}a^{-5n}\to 0.$
    As a consequence, for all large $n$,
    $\sqrt{1+|g_{n}^{\prime}(u)|^{2}}\leq 1+|g_{n}^{\prime}(u)|\leq 2+3^{n}a^{-4n}.$
    Hence
    ${\rm Vol}(A,\sigma)_{[-1,1]^{2}}\leq\log 3-4\log a.$
    It remains to prove the reverse inequality.
    For all large $n$, by Lemma 6.9 with $b=1$,
    $L_{n}\geq\int_{0}^{1}|g_{n}^{\prime}(u)|du\geq\int_{0}^{1}\Big|3^{n}a^{-4n}u^{3}\cos\!\left(\frac{a^{n}}{u}\right)\Big|du-1\geq 3^{n}a^{-4n}\cdot\frac{1}{2\pi(1+2\pi)^{4}}-1.$
    Hence
    ${\rm Vol}(A,\sigma)_{[-1,1]^{2}}\geq\log 3-4\log a.$
  • Case 3: $1<a\leq 3^{1/5}$.
    In this range, $3a^{-5}\geq 1$. For a lower bound, consider $0,u_{n}]$ with
    $u_{n}:=(3^{n}a^{-5n})^{-1/5}=3^{-n/5}a^{n}.$
    Then $|g_{n}(u)|\leq 1$ on $0,u_{n}]$, hence $0,u_{n}]\subset I_{n}$. Also note
    $\sqrt{1+|g_{n}^{\prime}(u)|^{2}}\geq|g_{n}^{\prime}(u)|\geq\left|3^{n}a^{-4n}u^{3}\cos\!\left(\frac{a^{n}}{u}\right)\right|-5\cdot 3^{n}a^{-5n}u^{4}.$
    Applying Lemma 6.9 with $b=u_{n}=3^{-n/5}a^{n}$, we have
    $\int_{0}^{u_{n}}\left|3^{n}a^{-4n}u^{3}\cos\!\left(\frac{a^{n}}{u}\right)\right|du\geq\frac{1}{2\pi\left(1+2\pi\right)^{4}}\cdot 3^{n/5}.$
    Hence
    $L_{n}\geq\int_{0}^{u_{n}}\left|3^{n}a^{-4n}u^{3}\cos\!\left(\frac{a^{n}}{u}\right)\right|du-\int_{0}^{u_{n}}5\cdot 3^{n}a^{-5n}u^{4}du\geq\frac{1}{2\pi\left(1+2\pi\right)^{4}}\cdot 3^{n/5}-1$
    which gives
    ${\rm Vol}(A,\sigma)_{[-1,1]^{2}}\geq\frac{1}{5}\log 3.$
    It remains to prove the reverse inequality. On $0,u_{n}]$,
    $\sqrt{1+|g_{n}^{\prime}(u)|^{2}}\leq|g_{n}^{\prime}(u)|+1\leq 3^{n}a^{-4n}u^{3}+5\cdot 3^{n}a^{-5n}u^{4}+1.$
    By a direct computation,
    $\limsup_{n\to+\infty}\frac{1}{n}\log\int_{0}^{u_{n}}\left(3^{n}a^{-4n}u^{3}+5\cdot 3^{n}a^{-5n}u^{4}+1\right)du\leq\frac{1}{5}\log 3.$
    So we only have to focus on the volume growth on the interval $u_{n},1]$.
    We next estimate the number of intervals of monotonicity for $g_{n}$ on $u_{n},1]$. Note that the critical points (i.e., $g^{\prime}_{n}(u)=0$) are determined by
    $5u\sin\!\frac{a^{n}}{u}=a^{n}\cos\!\frac{a^{n}}{u}.$
    Introducing the variable $t=a^{n}/u$, this equation is equivalent to
    $\tan t=\frac{t}{5}.$
    For $u\in[u_{n},1]$, we have
    $t\in\Bigl[a^{n},\frac{a^{n}}{u_{n}}\Bigr]=[a^{n},3^{n/5}].$
    On each interval $(k\pi-\frac{\pi}{2},k\pi+\frac{\pi}{2})$, the function $\tan t-t/5$ is strictly increasing, hence the above equation admits at most one solution. Therefore the number of solutions in $a^{n},3^{n/5}]$ is bounded by
    $\frac{3^{n/5}-a^{n}}{\pi}+1\leq\frac{3^{n/5}}{\pi}+1.$
    Each critical point contributes at most one change of monotonicity, so the number of intervals of monotonicity on $u_{n},1]$ is bounded by
    $\frac{3^{n/5}}{\pi}+2.$
    Restricted on $1,1]^{2}$, the length of the curve on each interval of monotonicity is at most $4$. Hence the volume growth rate on the interval $u_{n},1]$ restricted on $1,1]^{2}$ is bounded above by also $\frac{1}{5}\log 3.$ This proves the reverse inequality.
Lemma 6.9.
For any $a,b>0$ and $n\in\mathbb{N}$,
$\int_{0}^{b}u^{3}\left|\cos\frac{a^{n}}{u}\right|\,du\geq\frac{a^{4n}b^{4}}{2\pi\left(a^{n}+2\pi b\right)^{4}}.$
Proof.
By the change of variables $t=a^{n}/u$, we obtain
$\int_{0}^{b}u^{3}\left|\cos\frac{a^{n}}{u}\right|\,du=a^{4n}\int_{a^{n}/b}^{+\infty}t^{-5}|\cos t|\,dt.$
Let
$k_{0}:=\left\lceil\frac{a^{n}}{b\pi}\right\rceil.$
Since $t^{-5}|\cos t|\geq 0$, we have
$\int_{a^{n}/b}^{\infty}t^{-5}|\cos t|\,dt\geq\sum_{k\geq k_{0}}\int_{k\pi}^{(k+1)\pi}t^{-5}|\cos t|\,dt.$
For each $k\geq k_{0}$, using that $t^{-5}$ is decreasing on $k\pi,(k+1)\pi]$ and
$\int_{k\pi}^{(k+1)\pi}|\cos t|\,dt=2,$
we obtain
$\int_{k\pi}^{(k+1)\pi}t^{-5}|\cos t|\,dt\geq 2\,((k+1)\pi)^{-5}.$
Therefore,
$\int_{a^{n}/b}^{\infty}t^{-5}|\cos t|\,dt\geq 2\sum_{k\geq k_{0}}((k+1)\pi)^{-5}\geq\frac{2}{\pi}\int_{(k_{0}+1)\pi}^{\infty}t^{-5}\,dt=\frac{1}{2\pi}((k_{0}+1)\pi)^{-4}.$
Since $(k_{0}+1)\pi\leq a^{n}/b+2\pi$, it follows that
$\int_{a^{n}/b}^{\infty}t^{-5}|\cos t|\,dt\geq\frac{1}{2\pi}(a^{n}/b+2\pi)^{-4}.$
Consequently,
$\int_{0}^{b}u^{3}\left|\cos\frac{a^{n}}{u}\right|\,du\geq\frac{a^{4n}b^{4}}{2\pi\left(a^{n}+2\pi b\right)^{4}}.$
Acknowledgements
We are grateful to Mingyang Xia from Dalian University of Technology for many helpful and stimulating discussions during the early stages of this work. We also thank Jérôme Buzzi from Université Paris-Saclay for many valuable comments.
Yuntao Zang is supported by National Key R&D Program of China (2022YFA1005802) and NSFC (12201445, 12471186).